A Bibliography of Publications in Digital Technical Journal

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

17 August 2013 Version 1.19

Title word cross-reference

2 [Cou94]. μ [GBC^{+}92].

0.5 [MWA^{+}97]. 0.5-W [MWA^{+}97].

1 [CFF^{+}93, DLW94]. 100 [CM92, HKK86]. 160 [MWA^{+}97]. 160-MHz [MWA^{+}97].

2 [FG97]. 200-MHz [DWA^{+}92]. 2000 [PAFK95]. 21174 [Sch97].

3 [PP98]. 3-D [PP98]. 300 [BBB^{+}95]. 300-MHz [ERB^{+}95, BBB^{+}95]. 32 [HH97]. 32-b [MWA^{+}97]. 32-Bit [DSW86, FGJ^{+}88, Smi96]. 3500 [DeV88, Lid88]. 3500/3600 [DeV88, Lid88]. 3600 [DeV88, Lid88]. 3D [Lev94].

4 [GBC^{+}92]. 400 [BHH^{+}90, CCV90, DBB^{+}90, SCL^{+}90]. 4DT [LFB^{+}98].

5-series [ZML95]. 500/600 [BB93b].

600 [CBW92]. 6000 [Por92]. 6200 [All88b, All88a, BGHS88, MD88]. 64-bit [BBB^{+}95, DWA^{+}92, ERB^{+}95, Gok96, HS96, KSS96, NK98, Smi96].

7000 [AvI92]. 78032 [DSW86]. 78034 [FGJ^{+}88]. 78132 [BFL^{+}86].

8600 [BJ85, FGB85, FME85, Hac85, KE85, McE85, TCQ^{+}85]. 8800 [Mis87]. 8974 [PEWQ87].

90 [CM92]. 9000 [BH90, BBM^{+}90, CBB^{+}90, FFM90, GDL90].

ACA [Pat93]. Accelerator [SU96, Lev94]. Accelerators [PCCW91]. Access [BS98, JKP91, PKGV95, CJ90]. ACCESS.bus [Sic91]. ACMSxp [BCDW95]. Ada [BBM88, Con88]. Adapter [Jai91b, Kal91, Reg91, SW91]. Adapters [Bah96, CU96]. Adding [HB96, Smi96, Zar95]. Addressable [HS96]. Addressing [Hay93]. Advances [Ham98]. aGlance [Asc93b, Asc93a]. aided [BGHS88]. Alias [Rei98]. Alpha [BBB^{+}95, Sit92, BS98, BNA98, CFF^{+}93, CMS92, CGL98, Con92, CKF94, Cou94, ERB^{+}95, HB96, KN95a, KBC^{+}92, MBC^{+}94, NK98, TCS92, WH97]. Alpha-based [BS98]. AlphaServer [CD96, Dam96, DKM96, FFG^{+}95, GM95, Hay94, Her96, KSS96, Rus94, SHK^{+}96]. AlphaStation [ZML95]. AltaVista [AW97, Shi97, SDLT97]. among [PP98]. Analysis [Cre94, JH86, Jai91b, Kal91, KS89, Rei98, RHD98, WNR96, DH89, Jai91a, KSS96, MPJK89]. Analyzing [BS98, PR96]. Annotated [Bre98]. Application [PP93, Sna91, Wal98]. Application-level [Wal98]. Applications [HBT93, JR89, KN95b, New95, PR96, PP98, ZHO91, BGM88, FG89, MB90, RFW93]. Approach [BK86, HMPW93, Law94, Por92]. Arbitration [rCQO87]. Architectural [All88b, MHS90, Win93]. Architecture [Bah96, BET91, CU96, DTW95, FG97, HKK86, Her96, HB96, LC92, LFB^{+}98, New95, Reg91, Sic94, SHK^{+}96, TCQ^{+}85, AN90, McG90, PCS93, Sit92, Tra90, Duf87, JH86, LOP86]. Architectures [BS98]. Area [FY87, HKK86, MSK86, Swe91, Lea91]. Array [Sic94]. ASIC [BFG^{+}95]. Aspects [Bat89, ACS88]. Assessing [Tho93]. assurance [JBD^{+}92]. Assuring [GTH94]. Audio [Bah96, CU96, Pan93]. Automated [Tal95]. Automatic [IF98, Law94]. Availability [BBDW87, RR91, RSW89, Lea91]. Available [CW96]. Avoidance [Wal98]. AXP [CKF94, AvI92, CFF^{+}93, CMS92, CRB92, Con92, Cou94, DMR^{+}94, DEK^{+}92, GTH94, KBC^{+}92, MSW92, MBC^{+}94, Sit92, Tho93].

b [MWA^{+}97]. Back [GBC^{+}92]. Back-end [GBC^{+}92]. Backup [GBD96]. base [VW89]. based [BS98, CHK98, DD93, Dav91, Gre88, KN95b]. Basic [Ham98]. benchmark [PSS96]. between [CM90]. Bibliography [Bre98]. Binary [HH97, SCK^{+}92, WNR96]. Bit [DSW86, BBB^{+}95, DWA^{+}92, ERB^{+}95, FGJ^{+}88, Gok96, HS96, KSS96, NK98, Smi96]. Bitonal [PCCW91]. Board [Mas86]. Box [FGB85, TCQ^{+}85]. bridge [BB93b]. bridge/router [BB93b]. Browsing [KW95]. Building [CG96, HL95, LBD^{+}96]. Bus [rCQO87, Gil88b, Sic91, Mas88]. Byte [HB96].

C

[IF98, Rei98, KLS94, PSS96, Rei98, RHD98]. Cache [DeV88]. Cached [SHK^{+}96]. CAD [DFF^{+}92, HE90, Hut86]. Capabilities [RBG^{+}96]. Capacity [Met92]. carrier [JBD^{+}92]. Cartridge [BS86]. Case [GTH94, GT93, MPPO86, YM93, CRB92, Pat93]. CDA [AO90, GJ90, Tra90]. Challenges [DFF^{+}92, LL98]. Changing [HMPW93]. CHANNEL [FG97, LBD^{+}96]. Channels [KS89]. Character [BB93a, Hae93, YM93]. Characterization [CHK98, CD96]. Chinese [Yau93]. Chip [And92, BFL^{+}86, DSW86, DFF^{+}92, FGJ^{+}88, KN95a, Mas88, MBC^{+}94, MWY88, Mor88, NAC^{+}94, TCS92, Win88, CS90, DBB^{+}90, NGBJ92]. Chips [SU96]. CI [rCQO87]. Circuit [BBB^{+}95]. Client [DTW95, Met92]. Client-Server [DTW95, Met92]. Clock [Dam96]. Clusters [HS98, LBD^{+}96]. CMCTL [Mor88]. CMOS [MWA^{+}97, BFG^{+}95, BBB^{+}95, DWA^{+}92, ERB^{+}95, GBC^{+}92, JBD^{+}92, Mor88, NGBJ92, ZFF92]. CMOS-4 [GBC^{+}92, NGBJ92]. Code [BB93a, Bre98, BNA98]. collaboration [CG96]. Collaborative [PP98]. Combining [HH97]. Commercial [Tal95]. commit [SJR91]. Common [LG91, RG96, CBS91, JKP91]. Communication [Duf87]. Compiler [BCD^{+}92, HAA^{+}98, SCH98, Web90, Con88]. Compilers [Rei98]. Compiling [HBB^{+}95]. Complexity [ZFF92]. Compliance [New95]. compound [Tra90]. Compression [BGU96, Cre94, Pan93]. Computer [BGHS88, Cou94]. Computer-aided [BGHS88]. computing [BS96, Cel96, MSW92]. Concept [KLSM87]. Concepts [BNA98]. Control [Sid89]. Controller [Bat89, LB89, Mor88, SMV91, Sch97]. Controllers [Sic94]. Cooling [KE85]. Copies [SCH98]. copy [Wal98]. Corporate [AW97]. Cost [Kno93, MBC^{+}94, Dam96, WH97]. Cost-focused [MBC^{+}94]. COTS [Cou94]. Coupled [LB89]. CPU [And92, BBB^{+}95, CS90, DFF^{+}92, Her96, KN95a, Mas86, MHS90]. Cray [KFW94]. Cross [MPPO86]. Cultural [Gre93]. Custom [Hut86, CDLL92]. CVAX [FGJ^{+}88, Mas88, MWY88, Mor88, Win88].

D [PP98]. DART [Wal98]. Data [Cre94, HGH91, Hay93, KW95, PKGV95, SCP^{+}91, Tho93, Wal98, YST91, Zar95]. Data-copy [Wal98]. Database [Gok96, JR89, RR91, CMS92]. Databases [YM93]. DB [PKGV95]. DBMS [RSW89]. DCE [PB97]. Debugger [LL98]. Debugging [Bre98, BNA98, PR96]. DEC [ZD88, AvI92, Asc93b, Asc93a, CFF^{+}93, DLW94, DMR^{+}94, DEK^{+}92, GT93, HL95, KLS94, MSW92, Rei98, RFJR93, SW91, Zar95]. DEC/Test [ZD88]. DECbridge [KHM91]. DECchip [MBC^{+}94, NAC^{+}94]. DECconcentrator [TKK91]. DECdecision [SJY90]. DECdta [BET91]. DECelms [Swe91]. DECmcc [SS93]. DECmodel [HK94]. DECnet [BK86, BFN93a, BFN93b, CRB92, DH89, FG89, FJS86, LC92, MMS86, MPPO86, MPJK89, RG96, Syl86, YMP^{+}93]. DECnet-DOS [MMS86]. DECnet-ULTRIX [FJS86]. DECnet-VAX [BK86]. DECnet/OSI [BFN93b, RG96, BFN93a, YMP^{+}93]. DECnet/SNA [MPPO86, MPJK89]. DECNIS [BB93b]. DECprint [LG91]. DECsafe [CW96]. DECspin [PP93]. DECstation [FNW90]. DECtalk [Hal96]. DECwindows [Gre90, Ken90, McG90, MB90, RV90]. DECwrite [CM90]. Defect [NM92]. Defining [HJZ93]. Definition [All88b]. Delivering [WNR96]. Demonstration [TCS92]. dense [NGBJ92]. departmental [MSW92]. Deployment [Tal95]. Description [AvI92, CW96]. Design [And92, BGHS88, BLLWW95, CGS96, CU96, CCK^{+}92, Dav91, DeV88, DFF^{+}92, DKM96, FFM90, FPA^{+}96, GR94, Gok96, Gol87, Hay94, Her96, LG91, MSW92, NE89, NK98, RFJR93, RBC92, RG96, Roz96, Sch97, Sic94, SMV91, Sna91, SHK^{+}96, SW91, SS93, Syl86, WS86, WH97, WBW96, ZML95, CS90, CBB^{+}90, CG93, CM92, DDMS90, DEK^{+}92, HE90, HK94, SFK^{+}92, SJY90]. Designing [BJ85, GBD96, LL98, SJR91]. Desktop [Asc93b, PP93, PP98, Sic91, Asc93a, CM92]. detection [BH90]. Develop [DBMI92]. Developing [Mas86, Shi97]. Development [BBM88, FFG^{+}95, FNW90, GBC^{+}92, Gil88a, Gra88, GJ90, HL95, HS98, HMPW93, HBT91, KHM91, MMO89, Mas88, MWY88, NAC^{+}94, TCS92, TC90, Tre97, Wel92, YST91, Bea88, BGM88, RV90, SJY90]. Device [Nou92, SFK^{+}92]. Differences [Gre93]. Different [PP98]. Differential [McK98]. Digital [BGU96, BET91, CG93, Hay93, Pan93, Roz96, Sid89, BNA98, HH97, IF98, Rei98, Sch97, Tre97, WH97, BFG^{+}95, BLLWW95, CGS96, CMS92, CM90, DD93, Har93, HBMT96, JH86, KN95b, LOP86, LT90, MBC^{+}94, NAC^{+}94, PR96, SS91, Swe91, YJ90]. dimensional [RFN90]. directed [Law94]. Director [SS93]. Directory [OHS96, RG96]. Disk [Cra89, KS89, Sid89]. Display [Ken90]. Distributable [DTW95]. Distributed [BCDW95, BET91, Gol87, HBB^{+}95, HGH91, KLSM87, SCP^{+}91, ST87, BGM88, FG89, HJZ93, MMO89, NE89, Win93]. Distributed-memory [HBB^{+}95]. Distribution [Dam96]. DLT2000 [Cre94]. Document [tHMv93, CM90, LT90, Tra90]. Domain [Smi89]. Drive [BS86, Cra89, Cre94]. Drivers [Nou92]. Drives [Sid89]. Dual [DWA^{+}92]. Dual-Issue [DWA^{+}92]. DXML [KHM94].

Editor [Ano94b, Bla95b, Bla96a, Bla96b, Bla96c, Bla96d, Bla97, CM90, Lup88, Ano98]. Editors [Ano91a, Ano91b, Ano91c, Ano91d, Ano92a, Ano92b, Ano92c, Ano93a, Ano93b, Ano93c, Ano93d, Ano94a, Ano94c, Ano94d, Ano95, Bla95a]. Effects [HB96]. Efforts [DD93]. Electromigration [CAL92]. Electronic [LPWH95]. Emulation [HH97, MK86]. end [FFG^{+}95, GBC^{+}92]. Engineering [BGHS88, AO90, Goo88]. enhancement [CDLL92]. Enrollment [Con92]. Environment [BNP96, CGS96, HL95, KN95b, Bea88, FG89, BLLWW95, CW96]. environments [GJ89]. error [Por92]. Ethernet [MB90, MSK86, Swe91]. Evaluation [KHRBED91, MD88, NAC^{+}94, FG89]. Evolution [CKF94, Hut86, MK86, PSS86, Mie89, SV90]. eXcursion [FPA^{+}96, GL92]. Executables [CGL98]. Expanders [Ham98]. experience [HBMT96]. Extend [AW97]. Extended [HKK86, Swe91]. Extending [HS96]. extension [Ken90].

F [FGB85]. Failover [HS98]. Family

[AL92, All88b, All88a, AvI92, Bah96, CU96]. Fast [GBD96, HS98, Wal98, NGBJ92]. Fault [EP89, BH90, BAM91]. fault-tolerant [BAM91]. FDDI [HBT91, Jai91a, Jai91b, Kal91, Reg91, Swe91, YST91]. features [CBB^{+}90, Con88, Lea91, MHS90]. Fiber [HGH91, SCP^{+}91]. File [BC92, Gol87, GBD96, HP96, JL96, RBC92, WBW96, NE89]. Film [Smi89]. Firewall [SDLT97]. First [MBC^{+}94, BAM91]. Floating [BFL^{+}86, FGB85, MWY88]. focused [MBC^{+}94]. Foreword [Bla98, Bon95, Cab91, Car92, Gor94, Har96, HH91, Hol94, Hut93, Kem91, Lar98, Lau93, Mar96, Mor93, Nem96, Pat95, Pes93, Sit95, Sup92a, Sup92b, Uli96]. format [CM90, LT90, YJ90]. Fortran [Cel96, HBB^{+}95, KLS94, Tre97]. fourth [Las88]. fourth-generation [Las88]. Frame [RT93]. Framework [GR94, KN95b]. Framework-based [KN95b]. Functional [KN95a, CS90]. FUSE [HL95, Zar95]. FX [HH97].

Gabriel [Tal95]. Gateway

[MPPO86, BB93b, MPJK89]. GEM

[BCD^{+}92]. generation

[BBB^{+}95, FGJ^{+}88, GM95, Las88]. generator [Gra88]. GIGAswitch [SKÖ^{+}94]. GKS [ACS88]. Global [SCH98, HJZ93]. Graphical [HL95, Gra88]. Graphics [LFB^{+}98, SU96, Lev94, RFN90].

handling [Por92]. Hardware

[CU96, PCCW91]. HDSC [DDMS90].

Heads [Smi89]. Hierarchical [BH90, LB89]. High [BFG^{+}95, CFF^{+}93, DKM96, FFG^{+}95, Gil88b, Jai91b, Kal91, KHM94, LFB^{+}98, PAFK95, Reg91, RSW89, SKÖ^{+}94, TCS92, WH97, ADH^{+}90, DEK^{+}92, LBD^{+}96, UBB^{+}92, HBB^{+}95]. High-end [FFG^{+}95]. High-Performance [SKÖ^{+}94, BFG^{+}95, CFF^{+}93, KHM94, LFB^{+}98, PAFK95, Reg91, TCS92, WH97, ADH^{+}90, DEK^{+}92, LBD^{+}96, UBB^{+}92]. High-Speed [Jai91b, Gil88b, Kal91]. history [CRB92]. Host [Dav91]. Host-based [Dav91]. Hot [Ham98, JBD^{+}92]. HPF [LL98]. HS [Sic94]. HS-series [Sic94]. HSC [Bat89].

I/O [DKM96, PAFK95, Rus94]. II [Dav91, Hut86, Mas86]. illustrate [MHS90]. Image [PCCW91, tHMv93]. imaging [Lev94]. Implementation [BBB^{+}95, BNA98, BFN93b, DLW94, EP89, FG97, Gol87, Hal96, Her96, NM92, Reg91, SCP^{+}91, TCQ^{+}85, BFN93a, HBMT96, NE89, RFW93, YMP^{+}93]. Improvement [DD93]. Improvements [Cra89]. Improving [GTH94]. Increase [GTH94]. Independent [HBT93, Law94]. Information [Asc93b, Asc93a]. Input [HBT93]. Instantiation [IF98]. Instruction [MK86, SCH98]. Instructions [HB96, MHS90]. Integrated [BK86]. Integrating [HP96, KN95b, OHS96, PB97, SU96, GL92]. Integration [Asc93b, New95, Asc93a, CJ90, Pat93]. Integrator [PKGV95]. Integrity [Hac85]. Interapplication [CJ90]. interchange [CM90, LT90, YJ90]. Interconnect [Ham98, CAL92]. Interconnection [GBC^{+}92]. Interface [GR94, HGH91, Mas88, RG96, SCP^{+}91, Wel92, Gre90, SV90]. Interfaces [Roz96]. Interfacing [Gil88b]. Internal [ERB^{+}95]. International [Gre93, Win93]. Internationalization [RB93, YM93]. Internet [AW97, CG96, HBMT96, Shi97]. Introduction [Ano91a, Ano91b, Ano91c, Ano91d, Ano92a, Ano92b, Ano92c, Ano93a, Ano93b, Ano93c, Ano93d, Ano94a, Ano94b, Ano94c, Ano94d, Ano95, Ano98, Bla95a, Bla95b, Bla96a, Bla96b, Bla96c, Bla96d, Bla97, Mor98]. IP [CFF^{+}93, HS98]. isolation [BH90]. Issue [DWA^{+}92, ERB^{+}95, KN95a]. issues [Win93].

J300 [Bah96, CU96]. Japanese

[HBT93, Yau93].

KAP [KLS94]. Kernal [LJL91]. Kerr [Smi89]. Kit [NAC^{+}94]. Korean [Yau93]. Kubota [Lev94].

LAN [Hay93]. LANBridge [HKK86]. LANcontroller [SW91]. Language [Lup88, Cel96, Con88, Gre90, Las88]. Language-sensitive [Lup88]. Languages [Yau93]. Large [NK98, Tal95, KSS96]. Layer [HBT91]. layers [RFW93]. Level [PEWQ87, DeV88, GBC^{+}92, Wal98]. Library [KHM94, Smi96]. line [GBD96]. Link [YST91]. Local [FY87, HKK86, MSK86, Swe91, Lea91]. Lock [ST87]. Log [GBD96]. Log-structured [GBD96]. logic [NGBJ92]. Logical [And92]. Low [Dam96, WH97]. Low-cost [Dam96, WH97].

m [GBC^{+}92]. Magnetic [KS89, Smi89]. Mail [Shi97, RV90]. MAILbus [Mie89]. Making [SGZ86]. Management [Buß94, Gok96, Joh94, JR89, LJL91, PSS86, SS93, tHMv93, CMS92, Con92, GJ89, SDS93, ZD88]. Manager [EP89, ST87, ZD88]. ManageWORKS [GR94]. Managing [Swe91, Con92]. manufacture [DDMS90]. Manufacturing [Asc93b, NM92, Asc93a, CDLL92]. Margin [KS89]. Maximizing [HAA^{+}98]. Measured [HB96]. Measurement [DH89, RHD98, Tho93, PSS96]. measurements [DH88]. Mechanisms [RSW89]. Megadoc [tHMv93]. Memory [BS98, HS96, Her96, KFW94, Mor88, NK98, Sch97, HBB^{+}95, KSS96, NGBJ92, FG97, LBD^{+}96]. message [LBD^{+}96]. message-passing [LBD^{+}96]. Method [HBT93]. methodology [CDLL92]. Methods [DFF^{+}92]. MHz [BBB^{+}95, DWA^{+}92, ERB^{+}95, MWA^{+}97]. Microarchitecture [Mis87]. Microprocessor [DSW86, DWA^{+}92, ERB^{+}95, FGJ^{+}88, Gil88b, MWA^{+}97, NAC^{+}94, ZFF92]. microprocessors [UBB^{+}92]. Microscopy [Smi89]. Microsoft [Nou92]. MicroVAX [BFL^{+}86, CM92, DeV88, DSW86, Hut86, Lan86, Lid88, Mas86, MK86, SGZ86]. Middleware [PKGV95]. military [Cou94]. Mixed [BNP96]. Model [BHH^{+}90, CS90, CBW92, CCV90, CCK^{+}92, CM92, DBB^{+}90, RB93, SCL^{+}90, CCLN92, SFSW90]. Modeling [BBDW87, JH86, Kno93, Met92, MPJK89]. Modern [Cel96]. Modification [WNR96]. Module [GM95, Lid88, SHK^{+}96, Tre97, SCL^{+}90]. Monitor [Syl86]. Monitors [SS91]. Motion [NLU93]. MPP [KFW94]. MS [Abr89]. MS-DOS [Abr89]. multichip [DDMS90]. Multicomputer [CGS96]. Multifunction [Win88]. Multimedia [RFJR93]. Multiple [KN95a, OHS96]. Multiple-issue [KN95a]. Multiprocessing [DLW94, DKM96, GM88, Gil88b]. Multiprocessor [HAA^{+}98, Hay94, LB89, TCS92]. multiprotocol [BB93b, CG93]. Multivendor [New95].

name [MMO89]. Network [HKK86, LOP86, Law94, Nou92, PSS86, SDLT97, SDS93, Tal95, GJ89, RFN90]. Network-directed [Law94]. network-transparent [RFN90].

Networked [KW95, PP93]. Networking [BK86, CFF^{+}93, MPPO86, PAFK95, SCP^{+}91, Wal98, Har93, JH86]. Networks [AW97, HS98, MSK86, Swe91, RT93]. NMCC [Syl86]. NMCC/DECnet [Syl86]. Node [CBS91]. NOTES [Gil88a]. NT [CHK98, CGL98, FPA^{+}96]. NT-based [CHK98]. Numerical [SFK^{+}92]. NVAX [And92, DFF^{+}92, UBB^{+}92].

O [DKM96, PAFK95, Rus94]. Object [RFJR93, WNR96]. ObjectBroker [PB97]. Observations [Smi89]. on-chip [NGBJ92]. On-line [GBD96]. only [BGU96]. Open [BCDW95, DTW95, PKGV95, Sic91, Har93, RB93]. OpenVMS [BNP96, CRB92, GTH94, HS96, HP96, Joh94, KBC^{+}92, NK98, Tho93, Yau93, YMP^{+}93]. Operating [HP96, LJL91, Law94, Yau93]. Optimization [SCH98, WNR96, Web90]. Optimized [Bre98, BNA98, Wel92, SJR91]. Optimizing [BCD^{+}92, CGL98]. Option [Gok96]. Oracle7 [Gok96]. Ordering [Hae93]. Organization [ERB^{+}95]. OSF [CFF^{+}93, DLW94]. OSF/1 [CFF^{+}93, DLW94]. OSI [BFN93a, BFN93b, RFW93, RG96, YMP^{+}93]. Overview [AL92, All88a, CBS91, FME85, Har93, HGH91, JL96, KLSM87, LOP86, Lid88, Sto95, DBB^{+}90, McG90, Tra90].

Packaging [McE85]. Packet [SKÖ^{+}94]. Packet-switching [SKÖ^{+}94]. Parallel [Ham98, Cel96, BLLWW95]. Parallelizer [KLS94]. Parasight [PR96]. passing [LBD^{+}96]. past [BS96]. PATHWORKS [AL92, BC92, Met92, Nou92, RBC92, Wel92]. Patterns [BS98]. PC [CKF94]. PCI [NAC^{+}94]. Performance [Bat89, rCQO87, CD96, DMR^{+}94, DKM96, FG89, HAA^{+}98, HBB^{+}95, JH86, Jai91a, Jai91b, Kal91, KSS96, KHRBED91, MD88, PEWQ87, PSS96, RHD98, SKÖ^{+}94, ZFF92, ADH^{+}90, BFG^{+}95, CFF^{+}93, DEK^{+}92, KHM94, LBD^{+}96, LFB^{+}98, MSW92, MB90, PAFK95, Reg91, TCS92, UBB^{+}92, WH97]. Personal [Sch97, WH97]. PEX [RFN90]. Phase [Dav91]. PHIGS [ACS88]. Physical [Ham98, HBT91, BHH^{+}90]. Pictures [NLU93]. Pipelined [KN95a, TCQ^{+}85]. Platform [FFG^{+}95, KW95, Law94, SKÖ^{+}94, CMS92]. Platform-independent [Law94]. Playback [BGU96]. Plugging [Ham98]. Point [BFL^{+}86, FGB85, MWY88]. Pointer [BNP96, Smi96]. Policy [Buß94]. Port [DBMI92, GTH94]. Porting [CMS92, KBC^{+}92, Lan86]. PostScript [Ken90]. power [CBB^{+}90]. PowerStorm [LFB^{+}98]. practical [Win93]. Pragmatic [HMPW93, Por92]. Pragmatics [BBM88]. Preliminary [ZHO91]. Preserving [BS96]. Printer [JKP91, LG91]. PrintServer [SMV91]. Private [SDLT97]. Process [All88b, Asc93b, DD93, GBC^{+}92, GTH94, GT93, ZFF92, Asc93a, SFK^{+}92]. Processing [BCDW95, BET91, KHRBED91, PCCW91, RR91, SS91, ZHO91, BBM^{+}90, Gre88, SFSW90]. Processor [GM95, KE85, KN95a, Lid88, McE85, SHK^{+}96, CBW92, GDL90, SCL^{+}90]. Product [AL92, BK86, HMPW93, KHM91, MPPO86, TKK91, SJY90]. Production [JR89]. Productivity [GTH94, ACS88, BGM88, DH88]. Products [PSS86, ACS88, CMS92, DH89]. productvity [Con88]. Program [WNR96, Con92, Gra88]. Programmer [ACS88]. Programs [BS98, KLS94]. Project [Sto95, WS86, ZD88, Tal95]. Protecting [SDLT97]. protocol [SJR91, VW89, HBMT96, JKP91]. Provide [HS98]. provided [Con88].

Q22 [Mas88]. Q22-bus [Mas88]. QFD [HJZ93]. Quad [ERB^{+}95]. Quad-issue [ERB^{+}95]. qualification [CCV90]. Quality [GTH94, Kno93, Tho93].

RA90 [Cra89]. RALLY [Las88]. Rdb [DMR^{+}94, RFJR93]. Real [PR96]. Real-time [PR96]. Recognition [Roz96]. Recording [KS89]. Reduction [NM92]. Relational [Gok96, JR89, Las88]. relationship [CM90]. relay [RT93]. Reliability [BJ85, CAL92, JBD^{+}92]. Remote [GJ89, MB90]. Rendering [BGU96, SU96, Uli93, Lev94]. Repository [LPWH95]. requirements [HJZ93]. Research [KFW94]. Resolution [Buß94]. restoration [BS96]. revived [Cel96]. RISC [ERB^{+}95, MWA^{+}97, Web90]. Role [BGHS88]. Router [HS98, BB93b]. Routing [PCS93, CG93]. RQDX3 [WS86]. rule [Gre88]. rule-based [Gre88]. Rules [HMPW93]. Run [Smi96]. Run-time [Smi96].

sample [AN90]. scalar [SCL^{+}90]. SCAN [Gre88]. Scheduling [SCH98]. science [AO90]. Scientific [KHM94, Cel96]. SCSI [Ham98]. Second [BBB^{+}95, DeV88, FGJ^{+}88, GM95]. Second-generation [BBB^{+}95, FGJ^{+}88, GM95]. Second-level [DeV88]. Security [PB97]. SEI [DD93]. SEI-based [DD93]. Semantics [DTW95]. Semiconductor [ADH^{+}90]. sensitive [Lup88]. Sequoia [LPWH95, PAFK95, Sto95]. series [ZML95, FFG^{+}95, Sic94]. Server [BC92, CW96, DTW95, FFG^{+}95, Hay94, LB89, Met92, RBG^{+}96, RBC92, WBW96, AN90, MSW92]. Servers [MSK86]. service [GDL90, MMO89, NE89]. Services [OHS96, Abr89, FG89, Pat93]. set [CS90, DBB^{+}90]. Sets [NAC^{+}94]. Shadowing [Dav91, GTH94]. Shared [KFW94, PP98]. Sharing [PP98]. Signal [Hac85]. Simulation [DBMI92, BS96, SFK^{+}92]. Size [BNP96]. Sizing [ZHO91]. SNA [MPPO86, MPJK89]. Software [Bah96, BGU96, BLLWW95, CBS91, Con88, DBMI92, DH88, FJS86, Goo88, Gre93, GT93, Hal96, HL95, Kno93, KW95, Lan86, Law94, LFB^{+}98, McK98, NLU93, PAFK95, RSW89, Roz96, Shi97, Tal95, TCS92, Tho93, Bea88, BGM88, CG96, CG93, Gre88, ZD88]. Software-only [BGU96]. Sources [KW95]. Speech [Hal96, Roz96]. Speed [Jai91b, Gil88b, Kal91]. Spike [CGL98]. Spiralog [HP96, JL96, WBW96]. SQL [RBG^{+}96]. standard [Cou94]. Standards [New95]. Station [Joh94]. Storage [LB89]. StorageWorks [Sic94]. Strategies [NM92]. strategy [BH90, FFM90]. Strings [Hae93]. Structure [Joh94]. structured [GBD96]. Study [GTH94, GT93, MPPO86, YM93]. style [SV90]. Subjective [Tho93]. Subroutine [KHM94]. Subsystem [Rus94, SCP^{+}91]. SUIF [HAA^{+}98]. Suite [Hut86]. Superscalar [KN95a, SCH98]. Supervisor [LG91]. Support [LJL91, RFJR93, Smi96, Win88]. Supporting [Yau93]. switching [SKÖ^{+}94]. Sybase [RBG^{+}96]. Symmetric [DLW94, DKM96, GM88]. Synthesis [HE90]. System [BGHS88, BCD^{+}92, BJ85, rCQO87, CGS96, CHK98, Dam96, DKM96, FGB85, FME85, Gok96, Gol87, GBD96, Hac85, HP96, Hut86, JR89, KLSM87, LJL91, LG91, Lan86, Law94, MMS86, PEWQ87, RBG^{+}96, Sna91, SKÖ^{+}94, Win88, Yau93, tHMv93, ADH^{+}90, BH90, BAM91, BBM^{+}90, CBB^{+}90, Cou94, CCV90, FFM90, Gil88a, GJ89, HE90, KSS96, Las88, LBD^{+}96, RFN90, Duf87, HP96, JL96, Ken90, WBW96]. Systems [All88a, Asc93b, BNA98, Buß94, DMR^{+}94, FY87, HBB^{+}95, Hay94, KHRBED91, MK86, MD88, PP98, PEWQ87, Sid89, SCH98, WH97, Win88, Asc93a, CM92, DEK^{+}92, GL92, Lea91, MSW92, PSS96, Web90, Win93, CCK^{+}92, GM95].

table [YJ90]. Tape [BS86, Cre94]. TCP [CFF^{+}93]. TCP/IP [CFF^{+}93]. Tecate [KW95]. Technical [AvI92, CW96]. Techniques [ZHO91, DH89]. Technologies [ZFF92, JBD^{+}92]. Technology [Cra89, GBC^{+}92, Hal96, Ham98, Sid89, ADH^{+}90, BHH^{+}90, Cou94, NGBJ92]. Template [IF98]. Terminal [MSK86]. Terminals [EP91]. Test [CCV90, ZD88]. Testing [McK98, New95, VW89]. Text [Hal96, Gre88]. Text-to-Speech [Hal96]. Thin [Smi89]. Thin-Film [Smi89]. Three [DTW95, RFN90]. three-dimensional [RFN90]. Three-tier [DTW95]. tier [DTW95]. Tightly [LB89]. time [PR96, Smi96]. TK50 [BS86]. tolerant [BAM91]. Tool [PP98, Zar95]. toolkit [GJ90, TC90]. Tools [Asc93b, DFF^{+}92, HL95, WNR96, ZHO91, Asc93a, BGM88, SFK^{+}92, VW89]. TP [GT93]. TPC [PSS96]. TPC-C [PSS96]. Tracing [CHK98]. Transaction [BCDW95, BET91, DTW95, KHRBED91, LJL91, RR91, SS91, ZHO91, SJR91]. Transistor [JBD^{+}92, SFK^{+}92]. Translation [HH97, SCK^{+}92]. transparent [RFN90]. Transport [LC92, Wel92]. Triple [GBC^{+}92]. Triple-level [GBC^{+}92]. TruCluster [CGS96, PSS96]. Tunnel [AW97]. Turbo [SMV91]. two [DEK^{+}92, GL92].

UDP [CFF^{+}93]. UDP/IP [CFF^{+}93]. UltraSCSI [Ham98]. ULTRIX [BFN93a, BFN93b, FJS86, Lan86, RBC92, SCP^{+}91]. Unicode [BB93a]. unique [CBB^{+}90]. Unit [TCS92, DDMS90, GDL90]. Universal [BB93a, Hae93]. UNIX [CGS96, HBMT96, HL95, PR96]. Upgrades [Law94]. upper [RFW93]. usability [Goo88]. use [Cou94]. Used [Hut86, HE90]. User [GR94, Roz96, Gre90, SV90]. Using [AW97, DBMI92, Smi89, Tho93, KSS96, Pat93].

VALU [BGM88]. VALU-software [BGM88]. VAX [Abr89, ADH^{+}90, All88b, All88a, ACS88, BH90, BHH^{+}90, BGHS88, Bea88, BK86, BBM88, BJ85, BAM91, BBM^{+}90, CS90, CBB^{+}90, CBW92, Con88, CCV90, CCK^{+}92, CM92, DBB^{+}90, FFM90, FGB85, FME85, FGJ^{+}88, Gil88a, Gil88b, GDL90, Gre88, Hac85, HE90, KE85, KBC^{+}92, Las88, MMO89, McE85, Mis87, MD88, MHS90, NE89, PEWQ87, Por92, RSW89, ST87, SCL^{+}90, TCQ^{+}85, UBB^{+}92, ZD88]. VAX/VMS [Abr89, Bea88, ST87]. VAXcluster [BBDW87, rCQO87, FY87, KLSM87, Lea91, Sna91]. VAXsimPLUS [EP89]. VAXstation [CCLN92]. VAXvector [SFSW90]. Vector [BBM^{+}90, SFSW90]. Vendor [MPPO86]. Verification [And92, BAM91, KN95a, ZML95, CS90]. Version [DLW94, DMR^{+}94, FPA^{+}96, HBMT96, YMP^{+}93, ZD88]. Very [NK98, KSS96]. via [Wal98]. Video [Bah96, BGU96, CU96, Hay93, SU96, Uli93]. Videoconferencing [PP93]. Virtual [HS96, Nou92]. Visual [Tre97]. Visualization [Zar95]. Visualizing [KW95]. VLM [RBG^{+}96]. VLSI [CAL92, CDLL92, GBC^{+}92, NM92]. VMS [Abr89, Bea88, BC92, Dav91, GM88, LJL91, LG91, RV90, ST87]. Volume [Dav91, GTH94, Lev94]. VTX [BGM88].

W [MWA^{+}97]. WAVE [VW89]. While [GTH94]. Window [PP98, EP91]. windowing [GL92]. Windows [CHK98, CGL98, FPA^{+}96, GL92, HK94, Nou92]. within [Tre97]. without [SCH98]. Wizard [Tre97]. Word [HB96]. Workcenter [GT93]. Workflow [Buß94]. Workloads [CHK98, Met92]. Workstation [SGZ86, ZML95]. workstations [DEK^{+}92, Sch97, WH97].

X [EP91, RB93, SV90]. X/Open [RB93]. X11 [AN90]. XDPS [Ken90]. XUI [TC90].

yield [CDLL92].

References

Abrahams:1989:VSM

[Abr89] Alan Abrahams. VAX/VMS services for MS-DOS. Digital Technical Journal of Digital Equipment Corporation, 1 (9):44–50, June 1989. CODEN DTJOEL. ISSN 0898-901X.

Axtell:1988:PPA

[ACS88] B. A. Axtell, W. H. Clifford, and J. S. Saltz. Programmer productivity aspects of the VAX GKS and VAX PHIGS products. Digital Technical Journal of Digital Equipment Corporation, 1(6):62– 70, February 1988. CODEN DTJOEL. ISSN 0898-901X.

Adiletta:1990:STH

[ADH^{+}90] M. J. Adiletta, R. L. Doucette, J. H. Hackenberg, D. H. Leuthold, and D. M. Litwinetz. Semiconductor technology in a high-performance VAX system. Digital Technical Journal of Digital Equipment Corporation, 2(4):43–60, Fall 1990. CODEN DTJOEL. ISSN 0898-901X.

Abrahams:1992:OPP

[AL92] Alan Abrahams and David A. Low. An overview of the PATHWORKS product family. Digital Technical Journal of Digital Equipment Corporation, 4(1):8–14 (or 8– 15??), Winter 1992. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v4n1/An_Overview_of_the_
PATHWORKS_01oct1992DTJ501P8.
ps;   http://www.digital.
com:80/info/DTJ501/DTJ501SC.
TXT.

Allison:1988:OVF

[All88a] Brian Allison. An overview of the VAX 6200 family of systems. Digital Technical Journal of Digital Equipment Corporation, 1(7):10–18, August 1988. CODEN DTJOEL. ISSN 0898-901X.

Allison:1988:ADP

[All88b] Brian R. Allison. The architectural definition process of the VAX 6200 family. Digital Technical Journal of Digital Equipment Corporation, 1(7): 19–27, August 1988. CODEN DTJOEL. ISSN 0898-901X.

Angebranndt:1990:SXS

[AN90] S. Angebranndt and T. D.

Newman. The sample X11 server architecture. Digital Technical Journal of Digital Equipment Corporation, 2(3): 16–23, Summer 1990. CODEN DTJOEL. ISSN 0898-901X.

Anderson:1992:LVN

[And92] Walker Anderson. Logical verification of the NVAX CPU chip design. Digital Technical Journal of Digital Equipment Corporation, 4(3):38– 46, Summer 1992. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v4n3/Logical_Verification_
of_the_NV_01jan1993DTJ703P8.
ps;   http://www.digital.
com:80/info/DTJ703/DTJ703SC.
TXT.

Anonymous:1991:EIa

[Ano91a] Anonymous. Editors introduction. Digital Technical Journal of Digital Equipment Corporation, 3(1):??, Winter 1991. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/tp-introduction.
txt.

Anonymous:1991:EIb

[Ano91b] Anonymous. Editors introduction. Digital Technical Journal of Digital Equipment Corporation, 3(2):??, Spring 1991. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/fddi-introduction.
txt.

Anonymous:1991:EIc

[Ano91c] Anonymous. Editors introduction. Digital Technical Journal of Digital Equipment Corporation, 3(3):??, Summer 1991. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/vc-introduction.
txt.

Anonymous:1991:EId

[Ano91d] Anonymous. Editors introduction. Digital Technical Journal of Digital Equipment Corporation, 3(4):??, Fall 1991. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/im-introduction.
txt.

Anonymous:1992:EIa

[Ano92a] Anonymous. Editors introduction. Digital Technical Journal of Digital Equipment Corporation, 4(1):??, Winter 1992. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/pw-introduction.
txt.

Anonymous:1992:EIc

[Ano92b] Anonymous. Editors introduction. Digital Technical Journal of Digital Equipment Corporation, 4(3):??, Summer 1992. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/nv-introduction.
txt.

Anonymous:1992:EId

[Ano92c] Anonymous. Editors introduction. Digital Technical Journal of Digital Equipment Corporation, 4(4):??, Fall 1992. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/axp-introduction.
txt.

Anonymous:1993:EIa

[Ano93a]  Anonymous.  Editors in-

troduction. Digital Technical Journal of Digital Equipment Corporation, 5(1):??, Winter 1993. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        nw-01-introduction.txt.

Anonymous:1993:EIb

[Ano93b] Anonymous. Editors introduction. Digital Technical Journal of Digital Equipment Corporation, 5(2):??, Spring 1993. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

                                        digital.com/pub/Digital/
                                        info/DTJ/mm-01-introduction.
                                        txt.

Anonymous:1993:EIc

[Ano93c] Anonymous. Editors introduction. Digital Technical Journal of Digital Equipment Corporation, 5(3):??, Summer 1993. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

                                        digital.com/pub/Digital/
                                        info/DTJ/I18N-intro.txt.

Anonymous:1993:EId

[Ano93d] Anonymous. Editors introduction. Digital Technical Journal of Digital Equipment Corporation, 5(4):??, Fall 1993. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

                                        digital.com/pub/Digital/
                                        info/DTJ/swp-introduction.
                                        txt.

Anonymous:1994:EIa

[Ano94a] Anonymous. Editors introduction. Digital Technical Journal of Digital Equipment Corporation, 6(1):??, Winter 1994. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/sm-introduction.
txt.

Anonymous:1994:EIb

[Ano94b] Anonymous. Editor’s introduction. Digital Technical Journal of Digital Equipment Corporation, 6(2):??, Spring 1994. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v6n2-intro.txt.

Anonymous:1994:EIc

[Ano94c]  Anonymous.  Editors in-

troduction. Digital Technical Journal of Digital Equipment Corporation, 6(3):??, Summer 1994. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v6n3-introduction.txt.

Anonymous:1994:EId

[Ano94d] Anonymous. Editors introduction. Digital Technical Journal of Digital Equipment Corporation, 6(4):??, Fall 1994. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v6n4int.txt.

Anonymous:1995:EI

[Ano95] Anonymous. Editors introduction. Digital Technical Journal of Digital Equipment Corporation, 7(2):??, Spring 1995. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

                                        digital.com/pub/Digital/
                                        info/DTJ/v7n2-00-intro.txt.

Anonymous:1998:EI

[Ano98] Anonymous. Editor’s introduction. Digital Technical Journal of Digital Equipment Corporation, 9(3):??, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

                                        /www.digital.com:80/info/
                                        DTJR00/v9n3-intro.htm.

Appel:1990:CSE

[AO90] Neal B. Appel and Ronald M. Olson. CDA in science and engineering. Digital Technical Journal of Digital Equipment Corporation, 2(1):83–89, Winter 1990. CODEN DTJOEL. ISSN 0898-901X.

Ascher:1993:DDT

[Asc93a] D. Ascher. DEC@aGlance — integration of desktop tools and manufacturing process information systems. Digital Technical Journal of Digital Equipment Corporation, 5(2):100– 112, Spring 1993. CODEN DTJOEL. ISSN 0898-901X.

Ascher:1993:DAI

[Asc93b] David Ascher. DEC @aGlance — integration of desktop tools and manufacturing process information systems. Digital Technical Journal of Digital Equipment Corporation, 5(2): ??, Spring 1993. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v5n2/DEC_aGlanceIntegration_
of_De_01oct1993DTJA09SC.
txt;   http://www.digital.
com:80/info/DTJA09/DTJA09SC.
TXT.

Allison:1992:TDD

[AvI92] Brian R. Allison and Catharine van Ingen. Technical description of the DEC 7000 and DEC 10000 AXP family. Digital Technical Journal of Digital Equipment Corporation, 4(4):100–110, Fall 1992. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v4n4/Technical_
Description_of_the_D_01apr1993DTJ806P8.
ps;   http://www.digital.
com:80/info/DTJ806/DTJ806SC.
TXT.

Alden:1997:ATU

[AW97] Kenneth F. Alden and Edward P. Wobber. The AltaVista tunnel: Using the Internet to extend corporate networks. Digital Technical Journal of Digital Equipment Corporation, 9(2):5–??, ???? 1997. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJQ01/DTJQ01AH.HTM;   http:
//www.digital.com:80/info/

                                        DTJQ01/DTJQ01HM.HTM;   http:
                                        //www.digital.com:80/info/
                                        DTJQ01/DTJQ01P8.PS;   http:
                                        //www.digital.com:80/info/
                                        DTJQ01/DTJQ01PF.PDF;   http:
                                        //www.digital.com:80/info/
                                        DTJQ01/DTJQ01SC.TXT.

Bahl:1996:JFV

[Bah96] Paramvir Bahl. The J300 family of video and audio adapters: Software architecture. Digital Technical Journal of Digital Equipment Corporation, 7(4): 34–51, March 1996. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v7n4/The_J300_Family_of_
                                        Video_20apr1996DTJK03P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJK03;   http:
                                        //www.digital.com:80/info/
                                        DTJK03/DTJK03AH.HTM;   http:
                                        //www.digital.com:80/info/
                                        DTJK03/DTJK03P8.PS;   http:
                                        //www.digital.com:80/info/
                                        DTJK03/DTJK03PF.PDF;   http:
                                        //www.digital.com:80/info/
                                        DTJK03/DTJK03SC.TXT.

Bruckert:1991:VFF

[BAM91] William F. Bruckert, Carlos Alonso, and James M. Melvin. Verification of the first faulttolerant VAX system. Digital Technical Journal of Digital Equipment Corporation, 3(1): 79–85, Winter 1991. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v3n1/Verification_of_the_
                                        First_Faul_01oct1991DTJ108P8.

ps;   http://www.digital.
com:80/info/DTJ108/DTJ108SC.
TXT.

Bates:1989:PAH

[Bat89] Kenneth H. Bates. Performance aspects of the HSC controller. Digital Technical Journal of Digital Equipment Corporation, 1(8):25–37, February 1989. CODEN DTJOEL. ISSN 0898-901X.

Bettels:1993:UUC

[BB93a] Jörgen Bettels and F. Avery Bishop. Unicode: A universal character code. Digital Technical Journal of Digital Equipment Corporation, 5(3): 21–31, Summer 1993. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v5n3/UNICODE_A_UNIVERSAL_
CHARACTE_01jan1994DTJB02P8.
ps;   http://www.digital.
com:80/info/DTJB02/DTJB02SC.
TXT.

Bryant:1993:DMB

[BB93b] Stewart F. Bryant and David L. A. Brash. The DECNIS 500/600 multiprotocol bridge/ router and gateway. Digital Technical Journal of Digital Equipment Corporation, 5 (1):84–98, Winter 1993. CO-DEN DTJOEL. ISSN 0898-

901X.  URL ftp://ftp.
digital.com/pub/Digital/
info/DTJ/v5n1/The_DECNIS_
500600_Multiprotoc_01jul1993DTJ907P8.
ps;   http://www.digital.

com:80/info/DTJ907/DTJ907SC.
TXT.

Bowhill:1995:CIS

[BBB^{+}95] William J. Bowhill, Shane L. Bell, Bradley J. Benschneider, Andrew J. Black, Sharon M. Britton, Ruben W. Castelino, Dale R. Donchin, John H. Edmondson, Harry R. Fair, III, Paul E. Gronowski, Anil K. Jain, Patricia L. Kroesen, Marc E. Lamere, Bruce J. Loughlin, Shekhar Mehta, Robert O. Mueller, Ronald P. Preston, Sribalan Santhanam, Timothy A. Shedd, Michael J. Smith, and Stephen C. Thierauf. Circuit implementation of a 300-MHz 64-bit second-generation CMOS Alpha CPU. Digital Technical Journal of Digital Equipment Corporation, 7(1):100– 118, Winter 1995. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v7n1/Circuit_Implementation_
of_a_30_01jul1995DTJH08P8.
ps;   http://www.digital.
com:80/info/DTJH08/DTJH08SC.
TXT.

Balkovich:1987:VAM

[BBDW87] Edward E. Balkovich, Prashant Bhabhalia, William R. Dunnington, and Thomas F. Weyant. VAXcluster availability modeling. Digital Technical Journal of Digital Equipment Corporation, 1(5):69–79, September 1987. CODEN DTJOEL. ISSN 0898-901X.

Brender:1988:PDV

[BBM88] R. F. Brender, B. R. Brett, and C. Z. Mitchell. Pragmatics in the development of VAX Ada. Digital Technical Journal of Digital Equipment Corporation, 1(6):91–100, February 1988. CODEN DTJOEL. ISSN 0898-901X.

Brunner:1990:VPV

[BBM^{+}90] R. A. Brunner, D. P. Bhandarkar, F. X. McKeen, B. Patel, W. J. Rogers, Jr., and G. L. Yoder. Vector processing on the VAX 9000 system. Digital Technical Journal of Digital Equipment Corporation, 2 (4):61–79, Fall 1990. CODEN DTJOEL. ISSN 0898-901X.

Bresnahan:1992:PVF

[BC92] Edward W. Bresnahan and Siu Yin Cheng. PATHWORKS for VMS file server. Digital Technical Journal of Digital Equipment Corporation, 4 (1):15–23 (or 15–22??), Winter 1992. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v4n1/PATHWORKS_
for_VMS_File_Server_01oct1992DTJ502P8.
ps;   http://www.digital.
com:80/info/DTJ502/DTJ502SC.
TXT.

Blickstein:1992:GOC

[BCD^{+}92] David S. Blickstein, Peter W.

Craig, Caroline S. Davidson, R. Neil Faiman, Jr., Kent D. Glossop, Richard B. Grove, Steven O. Hobbs, and William B. Noyce. The GEM optimizing compiler system. Digital Technical Journal of Digital Equipment Corporation, 4(4):121– 136, Fall 1992. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v4n4/The_GEM_Optimizing_
Compiler_Sy_01apr1993DTJ808P8.
ps;   http://www.digital.
com:80/info/DTJ808/DTJ808SC.
TXT.

Baafi:1995:AOD

[BCDW95] Robert K. Baafi, J. Ian Carrie, William B. Drury, and Oren L. Wiesler. ACMSxp open distributed transaction processing. Digital Technical Journal of Digital Equipment Corporation, 7(1):23– 33, Winter 1995. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v7n1/ACMSxp_Open_Distributed_
Transa_01jul1995DTJH02P8.
ps;   http://www.digital.
com:80/info/DTJH02/DTJH02SC.
TXT.

Beander:1988:VVS

[Bea88] B. Beander. VAX/VMS software development environment. Digital Technical Journal of Digital Equipment Corporation, 1(6):10–19, February 1988. CODEN DTJOEL. ISSN 0898-901X.

Bernstein:1991:DDD

[BET91] Philip A. Bernstein, William T.

Emberton, and Vijay Treba. DECdta — digital’s distributed transaction processing architecture. Digital Technical Journal of Digital Equipment Corporation, 3(1):10– 17, Winter 1991. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v3n1/DECdta_Digitals_Distribute_
01oct1991DTJ101P8.ps;   http:
//www.digital.com:80/info/
DTJ101/DTJ101SC.TXT.

Basmaji:1995:DHC

[BFG^{+}95] Jean H. Basmaji, Kay R.

Fisher, Frank W. Gatulis, Herbert R. Kolk, and James F. Rosencrans. Digital’s highperformance CMOS ASIC. Digital Technical Journal of Digital Equipment Corporation, 7(1):66–76, Winter 1995. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v7n1/Digitals_Highperformance_
CMO_01jul1995DTJH05SC.txt;
 http://www.digital.com:
80/info/DTJH05/DTJH05SC.
TXT.

Bidermann:1986:MFP

[BFL^{+}86] William R. Bidermann, Amnon Fisher, Burton M. Leary, Robert J. Simcoe, and William R. Wheeler. The MicroVAX 78132 floating point chip. Digital Technical Journal of Digital Equipment Corporation, 1(2): 24–36, March 1986. CODEN DTJOEL. ISSN 0898-901X.

Buxton:1993:UID

[BFN93a] K. A. Buxton, E. J. Ferris, and A. K. Nash. The ULTRIX implementation of DECnet/ OSI. Digital Technical Journal of Digital Equipment Corporation, 5(1):34–43, Winter 1993. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

                                        digital.com/pub/Digital/
                                        info/DTJ/v5n1/The_ULTRIX_
                                        Implementation_of_D_01jul1993DTJ903P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJ903/DTJ903SC.
                                        TXT.

Buxton:1993:UIO

[BFN93b] Kim A. Buxton, Edward J. Ferris, and Andrew K. Nash. The ULTRIX implementation of DECnet/OSI. Digital Technical Journal of Digital Equipment Corporation, 5 (1):1–10 (??), Winter 1993. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

                                        digital.com/pub/Digital/
                                        info/DTJ/v5n1/The_ULTRIX_
                                        Implementation_of_D_01jul1993DTJ903P8.
                                        ps.

Basmaji:1988:CAE

[BGHS88] Jean H. Basmaji, Glenn P. Garvey, Masood Heydari, and Arthur L. Singer. The role of computer-aided engineering in the design of the VAX 6200 system. Digital Technical Journal of Digital Equipment Corporation, 1(7):47–56, August 1988. CODEN DTJOEL. ISSN 0898-901X.

Benson:1988:VVP

[BGM88] L. E. Benson, M. Gianatassio, and K. L. McKeen. VTX and VALU-software productivity tools for distributed applications development. Digital Technical Journal of Digital Equipment Corporation, 1(6): 80–90, February 1988. CODEN DTJOEL. ISSN 0898-901X.

Bahl:1996:SOC

[BGU96] Paramvir Bahl, Paul S. Gauthier, and Robert A. Ulichney. Software-only compression, rendering, and playback of digital video. Digital Technical Journal of Digital Equipment Corporation, 7(4):52– 75, March 1996. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v7n4/Softwareonly_Compression_
Ren_20apr1996DTJK04P8.ps;
 http://www.digital.com:
80/info/DTJK04;   http://
www.digital.com:80/info/
DTJK04/DTJK04AH.HTM;   http:
//www.digital.com:80/info/
DTJK04/DTJK04P8.PS;   http:
//www.digital.com:80/info/
DTJK04/DTJK04PF.PDF;   http:
//www.digital.com:80/info/
DTJK04/DTJK04SC.TXT.

Barnard:1990:HFD

[BH90] K. E. Barnard and R. P.

Harokopus. Hierarchical fault detection and isolation strategy for the VAX 9000 system. Digital Technical Journal of Digital Equipment Corporation, 2(4):

130–141, Fall 1990. CODEN DTJOEL. ISSN 0898-901X.

Bartoszek:1990:VMP

[BHH^{+}90] John T. Bartoszek, Robert J. Hannemann, Stephen P. Hansen, Robert J. McCarty, and John C. Sweeney. VAX 6000 Model 400 physical technology. Digital Technical Journal of Digital Equipment Corporation, 2(2):52–63, Spring 1990. CODEN DTJOEL. ISSN 0898-901X.

Bruckert:1985:DRV

[BJ85] William B. Bruckert and Ronald E. Josephson. Designing reliability into the VAX 8600 system. Digital Technical Journal of Digital Equipment Corporation, 1(1):71–77, August 1985. CODEN DTJOEL. ISSN 0898-901X.

Beck:1986:DVP

[BK86] P. Beck and J. Krycka. The DECnet-VAX product — an integrated approach to networking. Digital Technical Journal of Digital Equipment Corporation, 1(3):88–99, September 1986. CODEN DTJOEL. ISSN 0898-901X.

Blake:1995:EIa

[Bla95a] Jane C. Blake. Editors introduction. Digital Technical Journal of Digital Equipment Corporation, 7(1):3– 4, Winter 1995. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.

com/pub/Digital/info/DTJ/
v7n1-00-intro.txt.

Blake:1995:EIb

[Bla95b] Jane C. Blake. Editor’s introduction. Digital Technical Journal of Digital Equipment Corporation, 7(3):2, Fall 1995. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ000/v7n1-00-intro.
pdf.

Blake:1996:EIa

[Bla96a] Jane C. Blake. Editor’s introduction. Digital Technical Journal of Digital Equipment Corporation, 7(4):??, March 1996. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJK00/v7n4-intro.htm.

Blake:1996:EIb

[Bla96b] Jane C. Blake. Editor’s introduction. Digital Technical Journal of Digital Equipment Corporation, 8(1):??, May 1996. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com/.i/info/
DTJL00/v8n1-intro.htm.

Blake:1996:EIc

[Bla96c] Jane C. Blake. Editor’s introduction. Digital Technical Journal of Digital Equipment Corporation, 8(2):??, October 1996. CODEN DTJOEL. ISSN 0898-901X.

URL http://www.digital.
com:80/DTJM00/v8n2-intro.
htm.

Blake:1996:EId

[Bla96d] Jane C. Blake. Editor’s introduction. Digital Technical Journal of Digital Equipment Corporation, 8(3):??, ???? 1996. CODEN DTJOEL. ISSN 0898-901X. URL http://www.

                                        digital.com/info/DTJN00/
                                        v8n3-intro.htm.

Blake:1997:EI

[Bla97] Jane C. Blake. Editor’s introduction. Digital Technical Journal of Digital Equipment Corporation, 9(2):??, ???? 1997. CODEN DTJOEL. ISSN 0898-901X. URL http:/

                                        /www.digital.com:80/info/
                                        DTJQ00/v9n2-intro.htm.

Blake:1998:F

[Bla98] William C. Blake. Foreword. Digital Technical Journal of Digital Equipment Corporation, 10(1):??, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

                                        /www.digital.com:80/info/
                                        DTJT00/v10n1-fore.htm.

Benson:1995:DDP

[BLLWW95] Edward G. Benson, David C. P. LaFrance-Linden, Richard A. Warren, and Santa Wiryaman. Design of Digital’s Parallel Software Environment. Digital Technical Journal of Digital Equipment Corporation, 7 (3):24–38, Fall 1995. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v7n3/Design_of_Digitals_

Parallel_S_02jan1996DTJJ02P8.
ps;   http://www.digital.
com:80/info/DTJJ02/DTJJ02AH.
HTM;   http://www.digital.
com:80/info/DTJJ02/DTJJ02P8.
PS;   http://www.digital.
com:80/info/DTJJ02/DTJJ02PF.
PDF;   http://www.digital.
com:80/info/DTJJ02/DTJJ02SC.
TXT.

Brender:1998:DOC

[BNA98] Ronald F. Brender, Jeffrey E.

Nelson, and Mark E. Arsenault. Debugging optimized code: Concepts and implementation on DIGITAL Alpha systems. Digital Technical Journal of Digital Equipment Corporation, 10(1):81–99, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJT07/DTJT07HM.HTM;   http:
//www.digital.com:80/info/
DTJT07/DTJT07P8.PS;   http:
//www.digital.com:80/info/
DTJT07/DTJT07PF.PDF;   http:
//www.digital.com:80/info/
DTJT07/DTJT07SC.TXT.

Benson:1996:OMP

[BNP96] Thomas R. Benson, Karen L.

Noel, and Richard E. Peterson. The OpenVMS mixed pointer size environment. Digital Technical Journal of Digital Equipment Corporation, 8(2):72–82, October 1996. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJM06/DTJM06AH.HTM;   http:
//www.digital.com:80/info/

                                        DTJM06/DTJM06HM.HTM;   http:
                                        //www.digital.com:80/info/
                                        DTJM06/DTJM06P8.PS;   http:
                                        //www.digital.com:80/info/
                                        DTJM06/DTJM06PF.PDF;   http:
                                        //www.digital.com:80/info/
                                        DTJM06/DTJM06SC.TXT.

Bonney:1995:F

[Bon95] Jean C. Bonney. Foreword. Digital Technical Journal of Digital Equipment Corporation, 7(3):3–4, Fall 1995. CO-DEN DTJOEL. ISSN 0898-

                                        901X.  URL ftp://ftp.
                                        digital.com/pub/Digital/
                                        info/DTJJ00/v7n3_01_02_intro_
                                        fore.txt.

Brender:1998:ABD

[Bre98]  Ronald F. Brender.  An

annotated bibliography on debugging optimized code. Digital Technical Journal of Digital Equipment Corporation, 10(1):??, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

                                        /www.digital.com:80/info/
                                        DTJT07/DTJT07BIHM.HTM.

Boone:1986:TCT

[BS86] Steven E. Boone and Guenter E. Schneider. The TK50 cartridge tape drive. Digital Technical Journal of Digital Equipment Corporation, 1(2): 86–98, March 1986. CODEN DTJOEL. ISSN 0898-901X.

Burnet:1996:PCP

[BS96] Maxwell M. Burnet and Robert M. Supnik. Preserving computing’s past: restoration and simulation. Digital Technical Journal of Digital Equipment Corporation, 8(3):23–38, December 1996. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v8n3/Preserving_
Computings_Past_R_07jan1997DTJN02P8.
ps;   http://artematrix.
org/archive/computer/restoration.
simulation.htm;   http://
www.digital.com:80/info/
DTJN02/DTJN02AH.HTM;   http:
//www.digital.com:80/info/
DTJN02/DTJN02HM.HTM;   http:
//www.digital.com:80/info/
DTJN02/DTJN02P8.PS;   http:
//www.digital.com:80/info/
DTJN02/DTJN02PF.PDF;   http:
//www.digital.com:80/info/
DTJN02/DTJN02SC.TXT.

Balle:1998:AMA

[BS98] Simon M. Balle and Susanne C.

Steely, Jr. Analyzing memory access patterns of programs on Alpha-based architectures. Digital Technical Journal of Digital Equipment Corporation, 9(4):21–32, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJS02/DTJS02AH.HTM;   http:
//www.digital.com:80/info/
DTJS02/DTJS02HM.HTM;   http:
//www.digital.com:80/info/
DTJS02/DTJS02P8.PS;   http:
//www.digital.com:80/info/
DTJS02/DTJS02PF.PDF;   http:
//www.digital.com:80/info/
DTJS02/DTJS02SC.TXT.

Bussler:1994:PRW

[Buß94] Christoph J. Bußler. Policy resolution in workflow management systems. Digital Technical Journal of Digital Equipment Corporation, 6(4): 26–49, Fall 1994. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v6n4/Policy_Resolution_in_
                                        Workflow_01apr1995DTJG02P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJG02/DTJG02SC.
                                        TXT.

Cabrinety:1991:F

[Cab91] Larry Cabrinety. Foreword. Digital Technical Journal of Digital Equipment Corporation, 3(4):??, Fall 1991. CO-DEN DTJOEL. ISSN 0898-

                                        901X.  URL ftp://ftp.
                                        digital.com/pub/Digital/
                                        info/DTJ/im-foreword.txt.

Clement:1992:ERV

[CAL92] J. Joseph Clement, Eugenia M. Atakov, and James R. Lloyd. Electromigration reliability of VLSI interconnect. Digital Technical Journal of Digital Equipment Corporation, 4(2): 114–125, Spring 1992. CODEN DTJOEL. ISSN 0898-901X.

Carchidi:1992:F

[Car92] Joseph A. Carchidi. Foreword. Digital Technical Journal of Digital Equipment Corporation, 4(1):??, Winter 1992. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/pw-foreword.txt.

Chin:1990:UFV

[CBB^{+}90] D. J. Chin, B. G. Brown, C. F.

Butala, L. L. Chang, S. J. Chenetz, G. E. Cotter, B. T. Lynch, T. Natarajan, and L. J. Salafia. The unique features of the VAX 9000 power system design. Digital Technical Journal of Digital Equipment Corporation, 2(4):102–117, Fall 1990. CODEN DTJOEL. ISSN 0898-901X.

Ciarfella:1991:OCN

[CBS91] Paul W. Ciarfella, David Benson, and David S. Sawyer. An overview of the Common Node software. Digital Technical Journal of Digital Equipment Corporation, 3(2):42– 52, Spring 1991. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
fddi-node-software.txt;
http://www.digital.com:80/
info/DTJ204/DTJ204SC.TXT.

Chisvin:1992:VMP

[CBW92] Lawrence Chisvin, Gregg A.

Bouchard, and Thomas M. Wenners. The VAX 6000 Model 600 processor. Digital Technical Journal of Digital Equipment Corporation, 4(3): 47–59, Summer 1992. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v4n3/The_VAX_6000_Model_
600_01jan1993DTJ704P8.ps;

                                         http://www.digital.com:
                                        80/info/DTJ704/DTJ704SC.
                                        TXT.

Crowell:1992:DVMa

[CCK^{+}92] Jonathan C. Crowell, Kwong-Tak A. Chui, Thomas E. Kopec, Samyojita A. Nadkarni, and Dean A. Sovie. Design of the VAX 4000 Model 400, 500, and 600 Systems. Digital Technical Journal of Digital Equipment Corporation, 4(3):60–72, Summer 1992. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

                                        digital.com/pub/Digital/
                                        info/DTJ/v4n3/Design_of_
                                        the_VAX_4000_01jan1993DTJ705P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJ705/DTJ705SC.
                                        TXT.

Callander:1992:VM

[CCLN92] Michael A. Callander, Sr., Lauren M. Carlson, Andrew R. Ladd, and Mitchell O. Norcross. The VAXstation 4000 Model 90. Digital Technical Journal of Digital Equipment Corporation, 4(3):82– 91, Summer 1992. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v4n3/The_VAXstation_4000_
                                        Model_90_01jan1993DTJ707P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJ707/DTJ707SC.
                                        TXT.

Croll:1990:TQV

[CCV90] John W. Croll, Larry T. Camilli, and Anthony J. Vaccaro. Test and qualification of the VAX 6000 Model 400 system. Digital Technical Journal of Digital Equipment Corporation, 2(2):73–83, Spring 1990. CODEN DTJOEL. ISSN 0898-901X.

Cvetanovic:1996:APC

[CD96] Zarka Cvetanovic and Darrel D. Donaldson. AlphaServer 4100 performance characterization. Digital Technical Journal of Digital Equipment Corporation, 8(4):3–??, ???? 1996. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJO01/DTJO01AH.HTM;   http:
//www.digital.com:80/info/
DTJO01/DTJO01HM.HTM;   http:
//www.digital.com:80/info/
DTJO01/DTJO01P8.PS;   http:
//www.digital.com:80/info/
DTJO01/DTJO01PF.PDF;   http:
//www.digital.com:80/info/
DTJO01/DTJO01SC.TXT.

Collica:1992:YEM

[CDLL92] Randall S. Collica, X. Joseph Dietrich, Rudolf Lambracht Jr., and D. G. Lau. A yield enhancement methodology for custom VLSI manufacturing. Digital Technical Journal of Digital Equipment Corporation, 4(2):83–99, Spring 1992. CODEN DTJOEL. ISSN 0898-901X.

Celmaster:1996:MFR

[Cel96] William N. Celmaster. Modern Fortran revived as the language of scientific parallel computing. Digital Technical Journal of Digital Equipment Corporation, 8(3):39– 45, December 1996. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v8n3/Modern_Fortran_Revived_
as_the_07jan1997DTJN03P8.
ps;   http://www.digital.
com:80/info/DTJN03/DTJN03AH.
HTM;   http://www.digital.
com:80/info/DTJN03/DTJN03HM.
HTM;   http://www.digital.
com:80/info/DTJN03/DTJN03P8.
PS;   http://www.digital.
com:80/info/DTJN03/DTJN03PF.
PDF;   http://www.digital.
com:80/info/DTJN03/DTJN03SC.
TXT.

Chang:1993:HPT

[CFF^{+}93] Chran-Ham Chang, Richard Flower, John Forecast, Heather Gray, William R. Hawe, Ashok P. Nadkarni, K. K. Ramakrishnan, Uttam N. Shikarpur, and Kathleen M. Wilde. High-performance TCP/IP and UDP/IP networking in DEC OSF/1 for Alpha AXP. Digital Technical Journal of Digital Equipment Corporation, 5(1):44– 61, Winter 1993. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v5n1/Highperformance_TCPIP_
and_UD_01jul1993DTJ904P8.
ps;   http://www.digital.
com:80/info/DTJ904/DTJ904SC.
TXT.

Cobb:1993:DMR

[CG93] Graham R. Cobb and Elliot C. Gerberg. Digital’s multiprotocol routing software design. Digital Technical Journal of Digital Equipment Corporation, 5(1):70– 83, Winter 1993. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v5n1/Digitals_Multiprotocol_
Routin_01jul1993DTJ906P8.
ps;   http://www.digital.
com:80/info/DTJ906/DTJ906SC.
TXT.

Chiu:1996:BCS

[CG96] Dah Ming Chiu and David M.

Griffin. Building collaboration software for the Internet. Digital Technical Journal of Digital Equipment Corporation, 8(3):66– 74, December 1996. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v8n3/Building_Collaboration_
Softwar_07jan1997DTJN06P8.
ps;   http://www.digital.
com:80/info/DTJN06/DTJN06AH.
HTM;   http://www.digital.
com:80/info/DTJN06/DTJN06HM.
HTM;   http://www.digital.
com:80/info/DTJN06/DTJN06P8.
PS;   http://www.digital.
com:80/info/DTJN06/DTJN06PF.
PDF;   http://www.digital.
com:80/info/DTJN06/DTJN06SC.
TXT.

Cohn:1998:OAE

[CGL98] R. S. Cohn, D. W. Goodwin, and P. G. Lowney. Optimizing Alpha executables on Windows NT with Spike. Digital Technical Journal of Digital Equipment Corporation, 9(4):3–??, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJS01/DTJS01AH.HTM;   http:
//www.digital.com:80/info/
DTJS01/DTJS01HM.HTM;   http:
//www.digital.com:80/info/
DTJS01/DTJS01P8.PS;   http:
//www.digital.com:80/info/
DTJS01/DTJS01PF.PDF;   http:
//www.digital.com:80/info/
DTJS01/DTJS01SC.TXT.

Cardoza:1996:DTM

[CGS96] Wayne M. Cardoza, Frederick S. Glover, and William E. Snaman, Jr. Design of the TruCluster multicomputer system for the Digital UNIX environment. Digital Technical Journal of Digital Equipment Corporation, 8(1):5–17, May 1996. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v8n1/Design_of_
the_TruCluster_Multi_04jun1996DTJL01P8.
ps;   http://www.digital.
com:80/info/DTJL01/DTJL01.
HTM;   http://www.digital.
com:80/info/DTJL01/DTJL01AH.
HTM;   http://www.digital.
com:80/info/DTJL01/DTJL01P8.
PS;   http://www.digital.
com:80/info/DTJL01/DTJL01PF.
PDF;   http://www.digital.

com:80/info/DTJL01/DTJL01SC.
TXT.

Casmira:1998:TCW

[CHK98] Jason P. Casmira, David P.

Hunter, and David R. Kaeli. Tracing and characterization of Windows NT-based system workloads. Digital Technical Journal of Digital Equipment Corporation, 10(1):6–21, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJT01/DTJT01HM.HTM;   http:
//www.digital.com:80/info/
DTJT01/DTJT01P8.PS;   http:
//www.digital.com:80/info/
DTJT01/DTJT01PF.PDF;   http:
//www.digital.com:80/info/
DTJT01/DTJT01SC.TXT.

Cheung:1990:IAI

[CJ90] Baldwin K. Cheung and Neal F. Jacobson. Interapplication access and integration. Digital Technical Journal of Digital Equipment Corporation, 2(1):49, Winter 1990. CODEN DTJOEL. ISSN 0898-901X.

Conroy:1994:EAA

[CKF94] David G. Conroy, Thomas E.

Kopec, and Joseph R. Falcone. The evolution of the Alpha AXP PC. Digital Technical Journal of Digital Equipment Corporation, 6(1):54– 65, Winter 1994. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v6n1/The_Evolution_of_the_

                                        Alpha_01jul1994DTJD04P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJD04/DTJD04SC.
                                        TXT.

Cohen:1990:RBD

[CM90] Seth S. Cohen and Wm. Eugene Morgan. The relationship between the DECwrite editor and the Digital document interchange format. Digital Technical Journal of Digital Equipment Corporation, 2 (1):73–82, Winter 1990. CO-DEN DTJOEL. ISSN 0898-901X.

Crowell:1992:DVMb

[CM92] Jonathan C. Crowell and David W. Maruska. The design of the VAX 4000 Model 100 and MicroVAX 3100 Model 90 desktop systems. Digital Technical Journal of Digital Equipment Corporation, 4(3): 73–81, Summer 1992. CO-DEN DTJOEL. ISSN 0898-

                                        901X.  URL ftp://ftp.
                                        digital.com/pub/Digital/
                                        info/DTJ/v4n3/The_Design_
                                        of_the_VAX_01jan1993DTJ706P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJ706/DTJ706SC.
                                        TXT.

Coffler:1992:PDD

[CMS92] Jeffrey A. Coffler, Zia Mohamed, and Peter M. Spiro. Porting Digital’s database management products to the Alpha AXP platform. Digital Technical Journal of Digital Equipment Corporation, 4(4): 153–164, Fall 1992. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v4n4/Porting_Digitals_Database_
Man_01apr1993DTJ810P8.ps;
http://www.digital.com:
80/info/DTJ810/DTJ810SC.
TXT.

Conti:1988:SPF

[Con88] R. A. Conti. Software productvity features provided by the Ada language and the VAX Ada compiler. Digital Technical Journal of Digital Equipment Corporation, 1(6): 51–61, February 1988. CODEN DTJOEL. ISSN 0898-901X.

Conklin:1992:EMM

[Con92] Peter F. Conklin. Enrollment management, managing the Alpha AXP program. Digital Technical Journal of Digital Equipment Corporation, 4(4):193–205 (or 193– 207?? or 192–208??), Fall 1992. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v4n4/Enrollment_
Management_Managin_01apr1993DTJ813P8.
ps;   http://www.digital.
com:80/info/DTJ813/DTJ813SC.
TXT.

Couranz:1994:SAA

[Cou94] Robert Couranz. The E^{2}COTS system and Alpha AXP technology: the new computer standard for military use. Digital Technical Journal of Digital Equipment Corporation, 6(2): 22–33, Spring 1994. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v6n2/The_E_01oct1994DTJE02P8.
ps;   http://www.digital.
com:80/info/DTJE02/DTJE02SC.
TXT.

Crane:1989:DDT

[Cra89] Barbara A. Crane. Disk drive technology improvements in the RA90. Digital Technical Journal of Digital Equipment Corporation, 1(8):46– 60, February 1989. CODEN DTJOEL. ISSN 0898-901X.

Colombo:1992:DOA

[CRB92] James V. Colombo, Pamela J. Rickard, and Paul Benoit. DECnet for OpenVMS AXP: a case history. Digital Technical Journal of Digital Equipment Corporation, 4(4):165– 180, Fall 1992. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v4n4/DECnet_for_OpenVMS_
AXP_A_01apr1993DTJ811P8.
ps;   http://www.digital.
com:80/info/DTJ811/DTJ811SC.
TXT.

Cressman:1994:ADC

[Cre94] David C. Cressman. Analysis of data compression in the DLT2000 tape drive. Digital Technical Journal of Digital Equipment Corporation, 6(2): 62–71, Spring 1994. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/

v6n2/Analysis_of_Data_Compression_
i_01oct1994DTJE05P8.ps;
http://www.digital.com:80/
info/DTJE05/DTJE05SC.TXT.

Calcagni:1990:VMC

[CS90] Richard E. Calcagni and Will Sherwood. VAX 6000 Model 400 CPU chip set functional design verification. Digital Technical Journal of Digital Equipment Corporation, 2(2): 64–72, Spring 1990. CODEN DTJOEL. ISSN 0898-901X.

Correll:1996:JFV

[CU96] Kenneth W. Correll and Robert A. Ulichney. The J300 family of video and audio adapters: Architecture and hardware design. Digital Technical Journal of Digital Equipment Corporation, 7(4): 20–33, March 1996. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v7n4/The_J300_Family_of_
Video_20apr1996DTJK02P8.
ps;   http://www.digital.
com:80/info/DTJK02;   http:
//www.digital.com:80/info/
DTJK02/DTJK02AH.HTM;   http:
//www.digital.com:80/info/
DTJK02/DTJK02P8.PS;   http:
//www.digital.com:80/info/
DTJK02/DTJK02PF.PDF;   http:
//www.digital.com:80/info/
DTJK02/DTJK02SC.TXT.

Cohen:1996:TDD

[CW96] Lawrence S. Cohen and John H.

Williams. Technical description of the DECsafe Available Server Environment. Digital Technical Journal of Digital Equipment Corporation, 7 (4):89–100, March 1996. CO-DEN DTJOEL. ISSN 0898-

901X.  URL ftp://ftp.
digital.com/pub/Digital/
info/DTJ/v7n4/Technical_
Description_of_the_D_20apr1996DTJK06P8.
ps;   http://www.digital.
com:80/info/DTJK06;   http:
//www.digital.com:80/info/
DTJK06/DTJK06AH.HTM;   http:
//www.digital.com:80/info/
DTJK06/DTJK06P8.PS;   http:
//www.digital.com:80/info/
DTJK06/DTJK06PF.PDF;   http:
//www.digital.com:80/info/
DTJK06/DTJK06SC.TXT.

Dame:1996:ALC

[Dam96] Roger A. Dame. The AlphaServer 4100 low-cost clock distribution system. Digital Technical Journal of Digital Equipment Corporation, 8(4):38–??, ???? 1996. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJO03/DTJO03AH.HTM;   http:
//www.digital.com:80/info/
DTJO03/DTJO03HM.HTM;   http:
//www.digital.com:80/info/
DTJO03/DTJO03P8.PS;   http:
//www.digital.com:80/info/
DTJO03/DTJO03PF.PDF;   http:
//www.digital.com:80/info/
DTJO03/DTJO03SC.TXT.

Davis:1991:DVV

[Dav91] Scott H. Davis. Design of VMS Volume Shadowing Phase II — host-based shadowing.

Digital Technical Journal of Digital Equipment Corporation, 3(3):7–15, Summer 1991. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v3n3/Design_of_
VMS_Volume_Shadowin_01apr1992DTJ301P8.
ps;   http://www.digital.
com:80/info/DTJ301/DTJ301SC.
TXT.

Durdan:1990:OVM

[DBB^{+}90] W. Hugh Durdan, William J.

Bowhill, John F. Brown, William V. Herrick, Richard C. Marcello, Sridhar Samudrala, G. Michael Uhler, and Nicholas Wade. An overview of the VAX 6000 Model 400 chip set. Digital Technical Journal of Digital Equipment Corporation, 2(2): 36–51, Spring 1990. CODEN DTJOEL. ISSN 0898-901X.

Darcy:1992:USD

[DBMI92] George A. Darcy III, Ronald F.

Brender, Stephen J. Morris, and Michael V. Iles. Using simulation to develop and port software. Digital Technical Journal of Digital Equipment Corporation, 4(4):181– 192, Fall 1992. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v4n4/Using_Simulation_to_
Develop_an_01apr1993DTJ812P8.
ps;   http://www.digital.
com:80/info/DTJ812/DTJ812SC.
TXT.

Davies:1993:SPI

[DD93] Neil L. M. Davies and Margaret M. Dumont. SEIbased process improvement efforts at Digital. Digital Technical Journal of Digital Equipment Corporation, 5(4): 59–68, Fall 1993. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v5n4/SEIbased_Process_Improvement_
                                        01apr1994DTJC05P8.ps;   http:
                                        //www.digital.com:80/info/
                                        DTJC05/DTJC05SC.TXT.

Dunbeck:1990:HMU

[DDMS90] P. B. Dunbeck, R. J. Dischler, J. B. McElroy, and F. J. Swiatowiec. HDSC and multichip unit design and manufacture. Digital Technical Journal of Digital Equipment Corporation, 2(4):80–89, Fall 1990. CODEN DTJOEL. ISSN 0898-901X.

Dutton:1992:DDA

[DEK^{+}92] Todd A. Dutton, Daniel Eiref, Hugh R. Kurth, James J. Reisert, and Robin L. Stewart. The design of the DEC 3000 AXP systems, two high-performance workstations. Digital Technical Journal of Digital Equipment Corporation, 4(4):66–81, Fall 1992. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

                                        digital.com/pub/Digital/
                                        info/DTJ/v4n4/The_Design_
                                        of_the_DEC_01apr1993DTJ804P8.
                                        ps;   http://www.digital.

com:80/info/DTJ804/DTJ804SC.
TXT.

DeVane:1988:DVS

[DeV88] Charles J. DeVane. Design of the MicroVAX 3500/3600 second-level cache. Digital Technical Journal of Digital Equipment Corporation, 1(7): 87–94, August 1988. CODEN DTJOEL. ISSN 0898-901X.

Donchin:1992:NCC

[DFF^{+}92] Dale R. Donchin, Timothy C.

Fischer, Thomas F. Fox, Victor Peng, Ronald P. Preston, and William R. Wheeler. The NVAX CPU chip: Design challenges, methods, and CAD tools. Digital Technical Journal of Digital Equipment Corporation, 4(3):24– 37, Summer 1992. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v4n3/The_NVAX_CPU_Chip_Design_
01jan1993DTJ702P8.ps;   http:
//www.digital.com:80/info/
DTJ702/DTJ702SC.TXT.

Duncan:1988:SPM

[DH88] A. Smith Duncan and T. J.

Harris. Software productivity measurements. Digital Technical Journal of Digital Equipment Corporation, 1(6): 20–27, February 1988. CODEN DTJOEL. ISSN 0898-901X.

DSilva:1989:MAT

[DH89] Vijay G. D’Silva and Ruei-Hsin Hsiao. Measurement and analysis techniques for DECnet products. Digital Technical Journal of Digital Equipment Corporation, 1(9):78–86, June 1989. CODEN DTJOEL. ISSN 0898-901X.

Duncan:1996:HPI

[DKM96] Samuel H. Duncan, Craig D. Keefer, and Thomas A. McLaughlin. High performance I/O design in the AlphaServer 4100 symmetric multiprocessing system. Digital Technical Journal of Digital Equipment Corporation, 8(4):61–??, ???? 1996. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJO05/DTJO05AH.HTM;   http:
//www.digital.com:80/info/
DTJO05/DTJO05HM.HTM;   http:
//www.digital.com:80/info/
DTJO05/DTJO05P8.PS;   http:
//www.digital.com:80/info/
DTJO05/DTJO05PF.PDF;   http:
//www.digital.com:80/info/
DTJO05/DTJO05SC.TXT.

Denham:1994:DOV

[DLW94] Jeffrey M. Denham, Paula Long, and James A. Woodward. DEC OSF/1 version 3.0 symmetric multiprocessing implementation. Digital Technical Journal of Digital Equipment Corporation, 6(3): 29–43, Summer 1994. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v6n3/DEC_OSF1_Symmetric_
Multiproce_01jan1995DTJF03P8.
ps;   http://www.digital.

com:80/info/DTJF03/DTJF03SC.
TXT.

Dimino:1994:PDR

[DMR^{+}94] Lucien A. Dimino, Rabah Mediouni, T. K. Rengarajan, Michael S. Rubino, and Peter M. Spiro. Performance of DEC Rdb version 6.0 on AXP systems. Digital Technical Journal of Digital Equipment Corporation, 6(1):23– 35, Winter 1994. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v6n1/Performance_of_DEC_
Rdb_Versio_01jul1994DTJD02P8.
ps;   http://www.digital.
com:80/info/DTJD02/DTJD02SC.
TXT.

Dobberpuhl:1986:MCM

[DSW86] Daniel W. Dobberpuhl, Robert M.

Supnik, and Richard T. Witek. The MicroVAX 78032 chip, A 32-bit microprocessor. Digital Technical Journal of Digital Equipment Corporation, 1(2): 12–23, March 1986. CODEN DTJOEL. ISSN 0898-901X.

Depledge:1995:ODT

[DTW95] Norman G. Depledge, William A.

Turner, and Alexandra Woog. An open, distributable, threetier client-server architecture with transaction semantics. Digital Technical Journal of Digital Equipment Corporation, 7(1):34– 42, Winter 1995. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.

com/pub/Digital/info/DTJ/
v7n1/An_Open_Distributable_
Three_01jul1995DTJH03P8.
ps;   http://www.digital.
com:80/info/DTJH03/DTJH03SC.
TXT.

Duffy:1987:SCA

[Duf87] Darrell J. Duffy. The System Communication Architecture. Digital Technical Journal of Digital Equipment Corporation, 1(5):22–28, September 1987. CODEN DTJOEL. ISSN 0898-901X.

Dobberpuhl:1992:DCM

[DWA^{+}92] Daniel W. Dobberpuhl, Richard T. Witek, Randy Allmon, Robert Anglin, David Bertucci, Sharon Britton, Linda Chao, Robert A. Conrad, Daniel E. Dever, Bruce Gieseke, Soha M. N. Hassoun, Gregory W. Hoeppner, Kathryn Kuchler, Maureen Ladd, Burton M. Leary, Liam Madden, Edward J. McLellan, Derrick R. Meyer, James Montanaro, Donald A. Priore, Vidya Rajagopalan, Sridhar Samudrala, and Sribalan Santhanam. A 200-MHz 64-bit dual-issue CMOS microprocessor. Digital Technical Journal of Digital Equipment Corporation, 4(4):35– 50, Fall 1992. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v4n4/A_200MHz_64bit_Dualissue_
CM_01apr1993DTJ802P8.ps;
 http://www.digital.com:

80/info/DTJ802/DTJ802SC.
TXT.

Emlich:1989:VFM

[EP89] Larry W. Emlich and Herman D. Polich. VAXsimPLUS, A fault manager implementation. Digital Technical Journal of Digital Equipment Corporation, 1(8):38–45, February 1989. CODEN DTJOEL. ISSN 0898-901X.

Engberg:1991:XWT

[EP91] Björn Engberg and Thomas Porcher. X Window terminals. Digital Technical Journal of Digital Equipment Corporation, 3(4):26– 35, Fall 1991. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v3n4/X_Window_Terminals_
01jul1992DTJ402P8.ps;   http:
//www.digital.com:80/info/
DTJ402/DTJ402SC.TXT.

Edmondson:1995:IOA

[ERB^{+}95] John H. Edmondson, Paul I.

Rubinfeld, Peter J. Bannon, Bradley J. Benschneider, Debra Bernstein, Ruben W. Castelino, Elizabeth M. Cooper, Daniel E. Dever, Dale R. Donchin Timothy C. Fischer, Anil K. Jain, Shekhar Mehta, Jeanne E. Meyer, Ronald P. Preston, Vidya Rajagopalan, Chandrasekhara Somanathan, Scott A. Taylor, and Gilbert M. Wolrich. Internal organization of the Alpha 21164, a 300-MHz 64-bit quad-issue CMOS RISC microprocessor. Digital Technical Journal of Digital Equipment Corporation, 7(1):119– 135, Winter 1995. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v7n1/Internal_Organization_
                                        of_the_A_01jul1995DTJH09P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJH09/DTJH09SC.
                                        TXT.

Fenwick:1995:ASH

[FFG^{+}95] David M. Fenwick, Denis J. Foley, William B. Gist, Stephen R. VanDoren, and Daniel Wissell. The AlphaServer 8000 Series: High-end server platform development. Digital Technical Journal of Digital Equipment Corporation, 7(1): 43–65, Winter 1995. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v7n1/The_AlphaServer_8000_
                                        Series_H_01jul1995DTJH04P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJH04/DTJH04SC.
                                        TXT.

Fite:1990:DSV

[FFM90] D. B. Fite, Jr., T. Fossum, and D. Manley. Design strategy for the VAX 9000 system. Digital Technical Journal of Digital Equipment Corporation, 2 (4):13–24, Fall 1990. CODEN DTJOEL. ISSN 0898-901X.

Finkelstein:1989:PED

[FG89] Eugene Finkelstein and Richard A. Grawin. Performance evaluation of distributed applications and services in the DECnet environment. Digital Technical Journal of Digital Equipment Corporation, 1(9):61–77, June 1989. CODEN DTJOEL. ISSN 0898-901X.

Fillo:1997:AIM

[FG97] M. Fillo and R. B. Gillett.

Architecture and implementation of MEMORY CHAN-NEL 2. Digital Technical Journal of Digital Equipment Corporation, 9(1):27–??, ???? 1997. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJP03/DTJP03AH.HTM;   http:
//www.digital.com:80/info/
DTJP03/DTJP03HM.HTM;   http:
//www.digital.com:80/info/
DTJP03/DTJP03P8.PS;   http:
//www.digital.com:80/info/
DTJP03/DTJP03PF.PDF;   http:
//www.digital.com:80/info/
DTJP03/DTJP03SC.TXT.

Fossum:1985:FBF

[FGB85] Tryggve Fossum, William R.

Grundmann, and Virginia C. Blaha. The F box, floating point in the VAX 8600 system. Digital Technical Journal of Digital Equipment Corporation, 1(1):43–53, August 1985. CODEN DTJOEL. ISSN 0898-901X.

Fox:1988:CVAX

[FGJ^{+}88] Thomas F. Fox, Paul E.

Gronowski, Anil K. Jain, Burton M. Leary, and Daniel G. Miner. The CVAX 78034 chip, a 32-bit second-generation VAX microprocessor. Digital Technical Journal of Digital Equipment Corporation, 1(7): 95–108, August 1988. CODEN DTJOEL. ISSN 0898-901X.

Forecast:1986:DUS

[FJS86] John Forecast, James L. Jackson, and Jeffrey A. Schriesheim. The DECnet-ULTRIX software. Digital Technical Journal of Digital Equipment Corporation, 1(3):100–107, September 1986. CODEN DTJOEL. ISSN 0898-901X.

Fossum:1985:OVS

[FME85] Tryggve Fossum, James B. McElroy, and William English. An overview of the VAX 8600 system. Digital Technical Journal of Digital Equipment Corporation, 1(1):8–23, August 1985. CODEN DTJOEL. ISSN 0898-901X.

Furlong:1990:DD

[FNW90] Thomas C. Furlong, Michael J. K. Nielsen, and Neil C. Wilhelm. Development of the DECstation 3100. Digital Technical Journal of Digital Equipment Corporation, 2(2): 84–88, Spring 1990. CODEN DTJOEL. ISSN 0898-901X.

Freitas:1996:DEV

[FPA^{+}96] John T. Freitas, James G. Peterson, Scot A. Aurenz, Charles P. Guldenschuh, and Paul J. Ranauro. Design of eXcursion version 2 for Windows, Windows NT, and Windows 95. Digital Technical Journal of Digital Equipment Corporation, 8(1):32– 45, May 1996. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v8n1/Design_of_eXcursion_
Version_2_04jun1996DTJL03P8.
ps;   http://www.digital.
com:80/info/DTJL03/DTJL03.
HTM;   http://www.digital.
com:80/info/DTJL03/DTJL03AH.
HTM;   http://www.digital.
com:80/info/DTJL03/DTJL03P8.
PS;   http://www.digital.
com:80/info/DTJL03/DTJL03PF.
PDF;   http://www.digital.
com:80/info/DTJL03/DTJL03SC.
TXT.

Fox:1987:LAV

[FY87] Michael S. Fox and John A.

Ywoskus. Local area VAXcluster systems. Digital Technical Journal of Digital Equipment Corporation, 1(5):56–68, September 1987. CODEN DTJOEL. ISSN 0898-901X.

Garver:1992:CBP

[GBC^{+}92] Marion M. Garver, Joseph M.

Bulger, Thomas E. Clark, J. H. Dubash, L. M. Ross, and D. J. Welch. CMOS-4 back-end process development for a VLSI 0.75 μ m triple-level interconnection technology. Digital Technical Journal of Digital Equipment Corporation, 4(2): 51–72, Spring 1992. CODEN DTJOEL. ISSN 0898-901X.

Green:1996:DFO

[GBD96] Russell J. Green, Alasdair C. Baird, and J. Christopher Davies. Designing a fast, on-line backup system for a log-structured file system. Digital Technical Journal of Digital Equipment Corporation, 8(2):32–45, October 1996. CODEN DTJOEL. ISSN 0898-901X. URL http:/

                                        /www.digital.com:80/info/
                                        DTJM03/DTJM03AH.HTM;   http:
                                        //www.digital.com:80/info/
                                        DTJM03/DTJM03HM.HTM;   http:
                                        //www.digital.com:80/info/
                                        DTJM03/DTJM03P8.PS;   http:
                                        //www.digital.com:80/info/
                                        DTJM03/DTJM03PF.PDF;   http:
                                        //www.digital.com:80/info/
                                        DTJM03/DTJM03SC.TXT.

Goldman:1990:VSP

[GDL90] M. S. Goldman, P. H. Dormitzer, and P. A. Leveille. The VAX 9000 service processor unit. Digital Technical Journal of Digital Equipment Corporation, 2(4):90–101, Fall 1990. CODEN DTJOEL. ISSN 0898-901X.

Gilbert:1988:DVN

[Gil88a] P. D. Gilbert. Development of the VAX NOTES system. Digital Technical Journal of Digital Equipment Corporation, 1(6): 117–124, February 1988. CO-DEN DTJOEL. ISSN 0898-901X.

Gillett:1988:IVM

[Gil88b] Richard B. Gillett, Jr. Interfacing a VAX microprocessor to a high-speed multiprocessing bus. Digital Technical Journal of Digital Equipment Corporation, 1(7):28–46, August 1988. CODEN DTJOEL. ISSN 0898-901X.

Griffin:1989:RSM

[GJ89] David M. Griffin and Brad C.

Johnson. Remote system management in network environments. Digital Technical Journal of Digital Equipment Corporation, 1(9):29–36, June 1989. CODEN DTJOEL. ISSN 0898-901X.

Gumbel:1990:DCT

[GJ90] Richard T. Gumbel and Martin L. Jack. Development of the CDA toolkit. Digital Technical Journal of Digital Equipment Corporation, 2(1):38–48, Winter 1990. CODEN DTJOEL. ISSN 0898-901X.

Giokas:1992:EWI

[GL92] Dennis G. Giokas and Andrew T. Leskowitz. eXcursion for Windows: integrating two windowing systems. Digital Technical Journal of Digital Equipment Corporation, 4(1): 56–67, Winter 1992. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
pw-excursion.txt;   http:
//www.digital.com:80/info/
DTJ507/DTJ507SC.TXT.

Gamache:1988:VSM

[GM88] Rodney N. Gamache and Kathleen D. Morse. VMS symmetric multiprocessing. Digital Technical Journal of Digital Equipment Corporation, 1(7): 57–63, February 1988. CODEN DTJOEL. ISSN 0898-901X.

Godiwala:1995:SPM

[GM95] Nitin D. Godiwala and Barry A. Maskas. The second-generation processor module for AlphaServer 2100 Systems. Digital Technical Journal of Digital Equipment Corporation, 7(1): 77–88, Winter 1995. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v7n1/The_Secondgeneration_
Processo_01jul1995DTJH06P8.
ps;   http://www.digital.
com:80/info/DTJH06/DTJH06SC.
TXT.

Gokhale:1996:DOO

[Gok96] Vipin V. Gokhale. Design of the 64-bit option for the Oracle7 relational database management system. Digital Technical Journal of Digital Equipment Corporation, 8(4):76–??, ???? 1996. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJO06/DTJO06AH.HTM;   http:
//www.digital.com:80/info/
DTJO06/DTJO06HM.HTM;   http:
//www.digital.com:80/info/
DTJO06/DTJO06P8.PS;   http:
//www.digital.com:80/info/
DTJO06/DTJO06PF.PDF;   http:

//www.digital.com:80/info/
DTJO06/DTJO06SC.TXT.

Goldstein:1987:DFS

[Gol87] Andrew C. Goldstein. The design and implementation of a distributed file system. Digital Technical Journal of Digital Equipment Corporation, 1(5): 45–55, September 1987. CO-DEN DTJOEL. ISSN 0898-901X.

Good:1988:SUE

[Goo88] M. D. Good. Software usability engineering. Digital Technical Journal of Digital Equipment Corporation, 1(6):125– 133, February 1988. CODEN DTJOEL. ISSN 0898-901X.

Gordon:1994:F

[Gor94] Scott A. Gordon. Foreword.

Digital Technical Journal of Digital Equipment Corporation, 6(2):??, Spring 1994. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v6n2-foreword.txt.

Giokas:1994:DMU

[GR94] Dennis G. Giokas and John C.

Rokicki. The design of ManageWORKS: A user interface framework. Digital Technical Journal of Digital Equipment Corporation, 6(4):63–74, Fall 1994. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v6n4/The_Design_
of_ManageWORKS_A_01apr1995DTJG04P8.
ps;   http://www.digital.

                                        com:80/info/DTJG04/DTJG04SC.
                                        TXT.

Grass:1988:DGP

[Gra88] S. J. Grass. Development of a graphical program generator. Digital Technical Journal of Digital Equipment Corporation, 1(6):101–109, February 1988. CODEN DTJOEL. ISSN 0898-901X.

Greenwood:1988:VSR

[Gre88] S. R. Greenwood. VAX SCAN: rule-based text processing software. Digital Technical Journal of Digital Equipment Corporation, 1(6):40–50, February 1988. CODEN DTJOEL. ISSN 0898-901X.

Greenwood:1990:DUI

[Gre90]  S. R. Greenwood.  The

DECwindows user interface language. Digital Technical Journal of Digital Equipment Corporation, 2(3):34–43, Summer 1990. CODEN DTJOEL. ISSN 0898-901X.

Greenwood:1993:ICD

[Gre93] Timothy G. Greenwood. International cultural differences in software. Digital Technical Journal of Digital Equipment Corporation, 5(3):8– 20, Summer 1993. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v5n3/INTERNATIONAL_CULTURAL_
                                        DIFFERE_01jan1994DTJB01P8.
                                        ps;   http://www.digital.

com:80/info/DTJB01/DTJB01SC.
TXT.

Guerrieri:1993:DTW

[GT93] Ernesto Guerrieri and Bruce J.

Taylor. DEC TP Workcenter: A software process case study. Digital Technical Journal of Digital Equipment Corporation, 5(4):47– 58, Fall 1993. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v5n4/DEC_TP_WORKcenter_A_
Software_01apr1994DTJC04P8.
ps;   http://www.digital.
com:80/info/DTJC04/DTJC04SC.
TXT.

Goleman:1994:IPI

[GTH94] William L. Goleman, Robert G.

Thomson, and Paul J. Houlihan. Improving process to increase productivity while assuring quality: A case study of the volume shadowing port to OpenVMS AXP. Digital Technical Journal of Digital Equipment Corporation, 6 (1):36–53, Winter 1994. CO-DEN DTJOEL. ISSN 0898-

901X.  URL ftp://ftp.
digital.com/pub/Digital/
info/DTJ/v6n1/Improving_
Process_to_Increase_01jul1994DTJD03P8.
ps;   http://www.digital.
com:80/info/DTJD03/DTJD03SC.
TXT.

Hall:1998:MMP

[HAA^{+}98] Mary W. Hall, Jennifer M.

Anderson, Saman P. Amarasinghe, Brian R. Murphy, Shih-Wei Liao, Eduoard Bugnion, and Monica S. Lam. Maximizing multiprocessor performance with the SUIF compiler. Digital Technical Journal of Digital Equipment Corporation, 10(1):71–80, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

                                        /www.digital.com:80/info/
                                        DTJT06/DTJT06HM.HTM;   http:
                                        //www.digital.com:80/info/
                                        DTJT06/DTJT06P8.PS;   http:
                                        //www.digital.com:80/info/
                                        DTJT06/DTJT06PF.PDF;   http:
                                        //www.digital.com:80/info/
                                        DTJT06/DTJT06SC.TXT.

Hackenberg:1985:SIV

[Hac85] John H. Hackenberg. Signal integrity in the VAX 8600 system. Digital Technical Journal of Digital Equipment Corporation, 1(1):61–65, August 1985. CODEN DTJOEL. ISSN 0898-901X.

Haentjens:1993:OUC

[Hae93]  René Haentjens.  The or-

dering of universal character strings. Digital Technical Journal of Digital Equipment Corporation, 5(3):43– 52, Summer 1993. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v5n3/THE_ORDERING_OF_UNIVERSAL_
                                        CHAR_01jan1994DTJB04P8.ps;
                                         http://www.digital.com:
                                        80/info/DTJB04/DTJB04SC.
                                        TXT.

Hallahan:1996:DST

[Hal96] William I. Hallahan. DECtalk software: Text-to-speech technology and implementation. Digital Technical Journal of Digital Equipment Corporation, 7(4):5–19, March 1996. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v7n4/DECtalk_Software_
TexttoSpee_20apr1996DTJK01P8.
ps;   http://www.digital.
com:80/info/DTJK01;   http:
//www.digital.com:80/info/
DTJK01/DTJK01AH.HTM;   http:
//www.digital.com:80/info/
DTJK01/DTJK01P8.PS;   http:
//www.digital.com:80/info/
DTJK01/DTJK01PF.PDF;   http:
//www.digital.com:80/info/
DTJK01/DTJK01SC.TXT.

Ham:1998:RAB

[Ham98] William E. Ham. Recent advances in basic physical technology for parallel SCSI: Ultra-SCSI, expanders, interconnect, and hot plugging. Digital Technical Journal of Digital Equipment Corporation, 9(3):6–??, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJR01/DTJR01AH.HTM;   http:
//www.digital.com:80/info/
DTJR01/DTJR01HM.HTM;   http:
//www.digital.com:80/info/
DTJR01/DTJR01P8.PS;   http:
//www.digital.com:80/info/
DTJR01/DTJR01PF.PDF;   http:
//www.digital.com:80/info/
DTJR01/DTJR01SC.TXT.

Harper:1993:ODO

[Har93] John Harper. Overview of Digital’s open networking. Digital Technical Journal of Digital Equipment Corporation, 5(1): 12–20, Winter 1993. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v5n1/Overview_of_Digitals_
                                        Open_Net_01jul1993DTJ901P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJ901/DTJ901SC.
                                        TXT.

Harbert:1996:F

[Har96] Don Harbert. Foreword. Digital Technical Journal of Digital Equipment Corporation, 8 (1):??, May 1996. CODEN DTJOEL. ISSN 0898-901X.

                                        URL http://www.digital.
                                        com/.i/info/DTJL00/v8n1-
                                        fore.htm.

Hayden:1993:LAD

[Hay93]  Peter C. Hayden.  LAN

addressing for digital video data. Digital Technical Journal of Digital Equipment Corporation, 5(2):77– 83, Spring 1993. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v5n2/LAN_Addressing_for_
                                        Digital_Vid_01oct1993DTJA07P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJA07/DTJA07SC.
                                        TXT.

Hayes:1994:DAM

[Hay94] Fidelma M. Hayes. Design of the AlphaServer multiprocessor server systems. Digital Technical Journal of Digital Equipment Corporation, 6 (3):8–19, Summer 1994. CO-DEN DTJOEL. ISSN 0898-

901X.  URL ftp://ftp.
digital.com/pub/Digital/
info/DTJ/v6n3/Design_of_
the_AlphaServer_Mult_01jan1995DTJF01P8.
ps;   http://www.digital.
com:80/info/DTJF01/DTJF01SC.
TXT.

Hunter:1996:MEA

[HB96] David P. Hunter and Eric B.

Betts. Measured effects of adding byte and word instructions to the Alpha architecture. Digital Technical Journal of Digital Equipment Corporation, 8(4):89–??, ???? 1996. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJO08/DTJO08AH.HTM;   http:
//www.digital.com:80/info/
DTJO08/DTJO08HM.HTM;   http:
//www.digital.com:80/info/
DTJO08/DTJO08P8.PS;   http:
//www.digital.com:80/info/
DTJO08/DTJO08PF.PDF;   http:
//www.digital.com:80/info/
DTJO08/DTJO08SC.TXT.

Harris:1995:CHP

[HBB^{+}95] Jonathan Harris, John A.

Bircsak, M. Regina Bolduc, Jill Ann Diewald, Israel Gale, Neil W. Johnson, Shin Lee, C. Alexander Nelson, and Carl D. Offner. Compiling High Performance Fortran for distributed-memory systems. Digital Technical Journal of Digital Equipment Corporation, 7(3):5–23, Fall 1995. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v7n3/Compiling_
High_Performance_For_02jan1996DTJJ01P8.
ps;   http://www.digital.
com:80/info/DTJJ01/DTJJ01AH.
HTM;   http://www.digital.
com:80/info/DTJJ01/DTJJ01P8.
PS;   http://www.digital.
com:80/info/DTJJ01/DTJJ01PF.
PDF;   http://www.digital.
com:80/info/DTJJ01/DTJJ01SC.
TXT.

Harrington:1996:IPV

[HBMT96] Daniel T. Harrington, James P. Bound, John J. McCann, and Matt Thomas. Internet Protocol version 6 and the Digital UNIX implementation experience. Digital Technical Journal of Digital Equipment Corporation, 8(3):5–22, December 1996. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v8n3/Internet_Protocol_Version_
6_an_07jan1997DTJN01P8.ps;
 http://www.digital.com:
80/info/DTJN01/DTJN01AH.
HTM;   http://www.digital.
com:80/info/DTJN01/DTJN01HM.
HTM;   http://www.digital.
com:80/info/DTJN01/DTJN01P8.
PS;   http://www.digital.
com:80/info/DTJN01/DTJN01PF.
PDF;   http://www.digital.
com:80/info/DTJN01/DTJN01SC.
TXT.

Hutchison:1991:DFP

[HBT91] Jerry D. Hutchison, Christopher Baldwin, and Bruce W. Thompson. Development of the FDDI physical layer. Digital Technical Journal of Digital Equipment Corporation, 3(2): 19–30, Spring 1991. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v3n2/Development_of_the_
FDDI_Physic_01jan1992DTJ202P8.
ps;   http://www.digital.
com:80/info/DTJ202/DTJ202SC.
TXT.

Honma:1993:JIM

[HBT93] Takahide Honma, Hiroyoshi Baba, and Kuniaki Takizawa. Japanese input method independent of applications. Digital Technical Journal of Digital Equipment Corporation, 5(3):97–107, Summer 1993. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v5n3/JAPANESE_INPUT_
METHOD_INDEPEND_01jan1994DTJB08P8.
ps;   http://www.digital.
com:80/info/DTJB08/DTJB08SC.
TXT.

Hooper:1990:SCS

[HE90] D. F. Hooper and J. C. Eck.

Synthesis in the CAD system used to design the VAX 9000 system. Digital Technical Journal of Digital Equipment Corporation, 2(4):118–129, Fall 1990. CODEN DTJOEL. ISSN 0898-901X.

Herdeg:1996:DIA

[Her96] Glenn A. Herdeg. Design and implementation of the AlphaServer 4100 CPU and memory architecture. Digital Technical Journal of Digital Equipment Corporation, 8(4):48–??, ???? 1996. CODEN DTJOEL. ISSN 0898-901X. URL http:/

                                        /www.digital.com:80/info/
                                        DTJO04/DTJO04AH.HTM;   http:
                                        //www.digital.com:80/info/
                                        DTJO04/DTJO04HM.HTM;   http:
                                        //www.digital.com:80/info/
                                        DTJO04/DTJO04P8.PS;   http:
                                        //www.digital.com:80/info/
                                        DTJO04/DTJO04PF.PDF;   http:
                                        //www.digital.com:80/info/
                                        DTJO04/DTJO04SC.TXT.

Hawe:1991:FDD

[HGH91] William R. Hawe, Richard Graham, and Peter C. Hayden. Fiber distributed data interface overview. Digital Technical Journal of Digital Equipment Corporation, 3(2):10–19 (or 1– 10??), Spring 1991. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v3n2/Fiber_Distributed_Data_
                                        Interfa_01jan1992DTJ201P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJ201/DTJ201SC.
                                        TXT.

Hayakawa:1991:F

[HH91] Howard H. Hayakawa and George S. Hoff. Foreword. Digital Technical Journal of Digital Equipment Corporation, 3(3):??, Summer 1991.

CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/vc-foreword.txt.

Hookway:1997:DFC

[HH97] R. J. Hookway and M. A.

Herdeg. DIGITAL FX!32: Combining emulation and binary translation. Digital Technical Journal of Digital Equipment Corporation, 9(1):3–??, ???? 1997. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJP01/DTJP01AH.HTM;   http:
//www.digital.com:80/info/
DTJP01/DTJP01HM.HTM;   http:
//www.digital.com:80/info/
DTJP01/DTJP01P8.PS;   http:
//www.digital.com:80/info/
DTJP01/DTJP01PF.PDF;   http:
//www.digital.com:80/info/
DTJP01/DTJP01SC.TXT.

Hrones:1993:DGR

[HJZ93] John A. Hrones, Jr., Benjamin C. Jedrey, Jr., and Driss Zaaf. Defining global requirements with distributed QFD. Digital Technical Journal of Digital Equipment Corporation, 5(4):36– 46, Fall 1993. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v5n4/Defining_Global_Requirements_
w_01apr1994DTJC03P8.ps;
http://www.digital.com:80/
info/DTJC03/DTJC03SC.TXT.

Hoover:1994:DDW

[HK94] Stewart V. Hoover and Gary L.

Kratkiewicz. The design of DECmodel for Windows. Digital Technical Journal of Digital Equipment Corporation, 6 (4):50–62, Fall 1994. CO-DEN DTJOEL. ISSN 0898-

901X.  URL ftp://ftp.
digital.com/pub/Digital/
info/DTJ/v6n4/The_Design_
of_DECmodel_for_01apr1995DTJG03P8.
ps;   http://www.digital.
com:80/info/DTJG03/DTJG03SC.
TXT.

Hawe:1986:ELA

[HKK86] William R. Hawe, Mark F. Kempf, and Alan J. Kirby. The extended local area network architecture and LANBridge

100. Digital Technical Journal of Digital Equipment Corporation, 1(3):54–72, September 1986. CODEN DTJOEL. ISSN 0898-901X.

Hart:1995:DFB

[HL95] Richard O. Hart and Glenn Lupton. DEC FUSE: Building a graphical software development environment from UNIX tools. Digital Technical Journal of Digital Equipment Corporation, 7(2):5– 19, Spring 1995. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v7n2/DEC_FUSE_Building_a_
Graphica_02oct1995DTJI01P8.
ps;   http://www.digital.
com:80/info/DTJI01/DTJI01AH.
HTM;   http://www.digital.
com:80/info/DTJI01/DTJI01P8.
PS;   http://www.digital.

com:80/info/DTJI01/DTJI01PF.
PDF;   http://www.digital.
com:80/info/DTJI01/DTJI01SC.
TXT.

Huntwork:1993:CRP

[HMPW93] Paul K. Huntwork, Douglas W.

Muzzey, Christine M. Pietras, and Dennis R. Wixon. Changing the rules: A pragmatic approach to product development. Digital Technical Journal of Digital Equipment Corporation, 5(4):18– 35, Fall 1993. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v5n4/Changing_the_Rules_
A_Pragmati_01apr1994DTJC02P8.
ps;   http://www.digital.
com:80/info/DTJC02/DTJC02SC.
TXT.

Holmes:1994:F

[Hol94] Steve Holmes. Foreword. Digital Technical Journal of Digital Equipment Corporation, 6 (3):??, Summer 1994. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v6n3-foreword.txt.

Howell:1996:ISF

[HP96] Mark A. Howell and Julian M. Palmer. Integrating the Spiralog File System into the OpenVMS operating system. Digital Technical Journal of Digital Equipment Corporation, 8(2):46–56, October 1996. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJM04/DTJM04AH.HTM;   http:
//www.digital.com:80/info/
DTJM04/DTJM04HM.HTM;   http:
//www.digital.com:80/info/
DTJM04/DTJM04P8.PS;   http:
//www.digital.com:80/info/
DTJM04/DTJM04PF.PDF;   http:
//www.digital.com:80/info/
DTJM04/DTJM04SC.TXT.

Harvey:1996:EOA

[HS96] Michael S. Harvey and Leonard S. Szubowicz. Extending Open-VMS for 64-bit addressable virtual memory. Digital Technical Journal of Digital Equipment Corporation, 8(2):57–71, October 1996. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJM05/DTJM05AH.HTM;   http:
//www.digital.com:80/info/
DTJM05/DTJM05HM.HTM;   http:
//www.digital.com:80/info/
DTJM05/DTJM05P8.PS;   http:
//www.digital.com:80/info/
DTJM05/DTJM05PF.PDF;   http:
//www.digital.com:80/info/
DTJM05/DTJM05SC.TXT.

Higginson:1998:DRC

[HS98] Peter L. Higginson and Michael C. Shand. Development of router clusters to provide fast failover in IP networks. Digital Technical Journal of Digital Equipment Corporation, 9(3):32–??, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJR02/DTJR02AH.HTM;   http:
//www.digital.com:80/info/
DTJR02/DTJR02HM.HTM;   http:

//www.digital.com:80/info/
DTJR02/DTJR02P8.PS;   http:
//www.digital.com:80/info/
DTJR02/DTJR02SC.TXT;   http:
//www.digital.com:80/info/
DTJR02/DTJR02PF.PDF.

Hutchings:1986:ECC

[Hut86] Anthony F. Hutchings. The evolution of the custom CAD suite used on the MicroVAX II system. Digital Technical Journal of Digital Equipment Corporation, 1(2):48–55, March 1986. CODEN DTJOEL. ISSN 0898-901X.

Hutchings:1993:F

[Hut93] Tony F. Hutchings. Foreword. Digital Technical Journal of Digital Equipment Corporation, 5(4):??, Fall 1993. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/swp-foreword.txt.

Itzkowitz:1998:ATI

[IF98] Avrum E. Itzkowitz and Lois D. Foltan. Automatic template instantiation in DIG-ITAL C++. Digital Technical Journal of Digital Equipment Corporation, 10(1):22–31, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJT02/DTJT02HM.HTM;   http:
//www.digital.com:80/info/
DTJT02/DTJT02P8.PS;   http:
//www.digital.com:80/info/
DTJT02/DTJT02PF.PDF;   http:
//www.digital.com:80/info/
DTJT02/DTJT02SC.TXT.

Jain:1991:PAF

[Jai91a] R. Jain. Performance analysis of FDDI. Digital Technical Journal of Digital Equipment Corporation, 3(3):78– 88, Summer 1991. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v3n3/Performance_Analysis_
                                        of_FDDIgh_01apr1992DTJ307P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJ307/DTJ307SC.
                                        TXT.

Jain:1991:PAH

[Jai91b] Raj Jain. Performance analysis of a high-speed FDDI adapter (or “performance analysis of FDDI”??). Digital Technical Journal of Digital Equipment Corporation, 3(3):1–11 (??), Summer 1991. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v3n3/Performance_Analysis_
                                        of_FDDIgh_01apr1992DTJ307P8.
                                        ps.

Jackson:1992:THC

[JBD^{+}92] Daniel B. Jackson, David A. Bell, Brian S. Doyle, B. J. Fishbein, and D. B. Krakauer. Transistor hot carrier reliability assurance in CMOS technologies. Digital Technical Journal of Digital Equipment Corporation, 4(2):100– 113, Spring 1992. CODEN DTJOEL. ISSN 0898-901X.

Jain:1986:PAM

[JH86] R. Jain and W. R. Hawe. Performance analysis and modeling of Digital’s Networking Architecture. Digital Technical Journal of Digital Equipment Corporation, 1(3):25–34, September 1986. CODEN DTJOEL. ISSN 0898-901X.

Jones:1991:CPA

[JKP91] James D. Jones, Ajay P.

Kachrani, and Thomas E. Powers. The Common Printer Access Protocol. Digital Technical Journal of Digital Equipment Corporation, 3(4): 55–60, Fall 1991. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v3n4/The_Common_Printer_
Access_Prot_01jul1992DTJ405P8.
ps;   http://www.digital.
com:80/info/DTJ405/DTJ405SC.
TXT.

Johnson:1996:OSF

[JL96] James E. Johnson and William A.

Laing. Overview of the Spiralog File System. Digital Technical Journal of Digital Equipment Corporation, 8(2):5–14, October 1996. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJM01/DTJM01AH.HTM;   http:
//www.digital.com:80/info/
DTJM01/DTJM01HM.HTM;   http:
//www.digital.com:80/info/
DTJM01/DTJM01P8.PS;   http:
//www.digital.com:80/info/
DTJM01/DTJM01PF.PDF;   http:

                                        //www.digital.com:80/info/
                                        DTJM01/DTJM01SC.TXT.

Johnson:1994:SOM

[Joh94] James E. Johnson. The structure of the OpenVMS Management Station. Digital Technical Journal of Digital Equipment Corporation, 6(4): 75–88, Fall 1994. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v6n4/The_Structure_of_the_
                                        OpenVMS_01apr1995DTJG05P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJG05/DTJG05SC.
                                        TXT.

Joshi:1989:RDM

[JR89] Ashok M. Joshi and Karen E. Rodwell. A relational database management system for production applications. Digital Technical Journal of Digital Equipment Corporation, 1(8): 99–109, February 1989. CO-DEN DTJOEL. ISSN 0898-901X.

Kalkunte:1991:PAH

[Kal91] Ramsesh S. Kalkunte. Performance analysis of a highspeed FDDI adapter. Digital Technical Journal of Digital Equipment Corporation, 3(3): 64–77, Summer 1991. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v3n3/Performance_Analysis_
                                        of_a_High_01apr1992DTJ306P8.
                                        ps;   http://www.digital.

com:80/info/DTJ306/DTJ306SC.
TXT.

Kronenberg:1992:POV

[KBC^{+}92] Nancy P. Kronenberg, Thomas R.

Benson, Wayne M. Cardoza, Ravindran Jagannathan, and Benjamin J. Thomas III. Porting OpenVMS from VAX to Alpha AXP. Digital Technical Journal of Digital Equipment Corporation, 4(4):111– 120, Fall 1992. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v4n4/Porting_OpenVMS_from_
VAX_to_01apr1993DTJ807P8.
ps;   http://www.digital.
com:80/info/DTJ807/DTJ807SC.
TXT.

Kalita:1985:CVP

[KE85] E. Brian Kalita and William English. Cooling the VAX 8600 processor. Digital Technical Journal of Digital Equipment Corporation, 1(1):66–70, August 1985. CODEN DTJOEL. ISSN 0898-901X.

Kempf:1991:F

[Kem91] Mark F. Kempf. Foreword.

Digital Technical Journal of Digital Equipment Corporation, 3(2):??, Spring 1991. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/fddi-foreword.txt.

Kent:1990:XDP

[Ken90] C. A. Kent. XDPS: a Display PostScript System extension for DECwindows. Digital Technical Journal of Digital Equipment Corporation, 2(3): 64–73, Summer 1990. CODEN DTJOEL. ISSN 0898-901X.

Koeninger:1994:SMM

[KFW94] R. Kent Koeninger, Mark Furtney, and Martin Walker. A shared memory MPP from Cray Research. Digital Technical Journal of Digital Equipment Corporation, 6(2):8– 21, Spring 1994. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v6n2/A_Shared_Memory_MPP_
from_01oct1994DTJE01P8.ps;
 http://www.digital.com:
80/info/DTJE01/DTJE01SC.
TXT.

Kochem:1991:DDP

[KHM91] Robert C. Kochem, James S. Hiscock, and Brian T. Mayo. Development of the DECbridge 500 product. Digital Technical Journal of Digital Equipment Corporation, 3(2):53–63 (or 53– 65??), Spring 1991. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v3n2/Development_of_the_
DECbridge_5_01jan1992DTJ204P8.
ps;   http://www.digital.
com:80/info/DTJ205/DTJ205SC.
TXT.

Kamath:1994:DHS

[KHM94] Chandrika Kamath, Roy Ho, and Dwight P. Manley. DXML: A high-performance scientific subroutine library. Digital Technical Journal of Digital Equipment Corporation, 6(3): 44–56, Summer 1994. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v6n3/DXML_A_Highperformance_
Scien_01jan1995DTJF04P8.
ps;   http://www.digital.
com:80/info/DTJF04/DTJF04SC.
TXT.

Kohler:1991:PET

[KHRBED91] Walter H. Kohler, Yun-Ping Hsu, Thomas K. Rogers, and Wael H. Bahaa-El-Din. Performance evaluation of transaction processing systems. Digital Technical Journal of Digital Equipment Corporation, 3(1):45–58 (or 45– 57??), Winter 1991. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v3n1/Performance_Evaluation_
of_Tran_01oct1991DTJ104P8.
ps;   http://www.digital.
com:80/info/DTJ104/DTJ104SC.
TXT.

Kuhn:1994:KPD

[KLS94] Robert H. Kuhn, Bruce Leasure, and Sanjiv M. Shah. The KAP parallelizer for DEC Fortran and DEC C programs. Digital Technical Journal of Digital Equipment Corporation, 6(3):57– 70, Summer 1994. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/

                                        v6n3/The_KAP_Parallelizer_
                                        for_DEC_01jan1995DTJF05P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJF05/DTJF05SC.
                                        TXT.

Kronenberg:1987:VC

[KLSM87] Nancy P. Kronenberg, Henry M. Levy, William D. Strecker, and Richard J. Merewood. The VAXcluster concept: An overview of a distributed system. Digital Technical Journal of Digital Equipment Corporation, 1(5):7–21, September 1987. CODEN DTJOEL. ISSN 0898-901X.

Kantrowitz:1995:FVM

[KN95a] Michael Kantrowitz and Lisa M. Noack. Functional verification of a multiple-issue, pipelined, superscalar Alpha processor

— the Alpha 21164 CPU chip. Digital Technical Journal of Digital Equipment Corporation, 7(1):136–144, Winter 1995. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

                                        digital.com/pub/Digital/
                                        info/DTJ/v7n1/Functional_
                                        Verification_of_a_M_01jul1995DTJH10P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJH10/DTJH10SC.
                                        TXT.

Kirkley:1995:IAD

[KN95b] James R. Kirkley and William G. Nichols. Integrating applications with Digital’s frameworkbased environment. Digital Technical Journal of Digital Equipment Corporation, 7(2): 47–55, Spring 1995. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v7n2/Integrating_Applications_
with_02oct1995DTJI04P8.ps;
http://www.digital.com:
80/info/DTJI04/DTJI04AH.
HTM;   http://www.digital.
com:80/info/DTJI04/DTJI04P8.
PS;   http://www.digital.
com:80/info/DTJI04/DTJI04PF.
PDF;   http://www.digital.
com:80/info/DTJI04/DTJI04SC.
TXT.

Knox:1993:MCS

[Kno93] Stephen T. Knox. Modeling the cost of software quality. Digital Technical Journal of Digital Equipment Corporation, 5 (4):9–17, Fall 1993. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v5n4/Modeling_the_Cost_of_
Software_01apr1994DTJC01P8.
ps;   http://www.digital.
com:80/info/DTJC01/DTJC01SC.
TXT.

Kretschmer:1989:MAM

[KS89] Reinhard Kretschmer and Siegbert Sadowski. Margin analysis on magnetic disk recording channels. Digital Technical Journal of Digital Equipment Corporation, 1(8):81– 87, February 1989. CODEN DTJOEL. ISSN 0898-901X.

Kawaf:1996:PAU

[KSS96] Tareef Kawaf, D. John Shakshober, and David C. Stanley. Performance analysis using very large memory on the 64-bit AlphaServer system. Digital Technical Journal of Digital Equipment Corporation, 8(3):58– 65, December 1996. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v8n3/Performance_Analysis_
Using_Ver_07jan1997DTJN05P8.
ps;   http://www.digital.
com:80/info/DTJN05/DTJN05AH.
HTM;   http://www.digital.
com:80/info/DTJN05/DTJN05HM.
HTM;   http://www.digital.
com:80/info/DTJN05/DTJN05P8.
PS;   http://www.digital.
com:80/info/DTJN05/DTJN05PF.
PDF;   http://www.digital.
com:80/info/DTJN05/DTJN05SC.
TXT.

Kochevar:1995:TSP

[KW95] Peter D. Kochevar and Leonard R. Wanger. Tecate: A software platform for browsing and visualizing data from networked data sources. Digital Technical Journal of Digital Equipment Corporation, 7 (3):66–83, Fall 1995. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v7n3/Tecate_A_Software_Platform_
fo_02jan1996DTJJ05P8.ps;
 http://www.digital.com:
80/info/DTJJ05/DTJJ05AH.
HTM;   http://www.digital.
com:80/info/DTJJ05/DTJJ05P8.
PS;   http://www.digital.
com:80/info/DTJJ05/DTJJ05PF.
PDF;   http://www.digital.

com:80/info/DTJJ05/DTJJ05SC.
TXT.

Lanza:1986:PUS

[Lan86] Raymond J. Lanza. Porting ULTRIX software to the MicroVAX system. Digital Technical Journal of Digital Equipment Corporation, 1(2): 99–105, March 1986. CODEN DTJOEL. ISSN 0898-901X.

Lary:1998:F

[Lar98] Richard Lary. Foreword. Digital Technical Journal of Digital Equipment Corporation, 9 (3):??, ???? 1998. CODEN DTJOEL. ISSN 0898-901X.

URL http://www.digital.
com:80/info/DTJR00/v9n3-
fore.htm.

Lasher:1988:VRS

[Las88] L. Lasher. The VAX RALLY system — a relational fourthgeneration language. Digital Technical Journal of Digital Equipment Corporation, 1(6): 71–79, February 1988. CODEN DTJOEL. ISSN 0898-901X.

Lauck:1993:F

[Lau93] Anthony G. Lauck. Foreword. Digital Technical Journal of Digital Equipment Corporation, 5(1):??, Winter 1993. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/nw-02-foreword.
txt.

Lawson:1994:ANO

[Law94] John R. Lawson, Jr. Automatic, network-directed operating system software upgrades: A platform-independent approach. Digital Technical Journal of Digital Equipment Corporation, 6(4):89–100, Fall 1994. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v6n4/Automatic_
Networkdirected_Op_01apr1995DTJG06P8.
ps;   http://www.digital.
com:80/info/DTJG06/DTJG06SC.
TXT.

Lary:1989:HSC

[LB89] Richard E. Lary and Robert G. Bean. The hierarchical storage controller, A tightly coupled multiprocessor as storage server. Digital Technical Journal of Digital Equipment Corporation, 1(8):8–24, February 1989. CODEN DTJOEL. ISSN 0898-901X.

Lawton:1996:BHM

[LBD^{+}96] James V. Lawton, John J. Brosnan, Morgan P. Doyle, Seosamh D. ÓRiordáin, and Timothy G. Reddin. Building a high-performance messagepassing system for MEMORY CHANNEL clusters. Digital Technical Journal of Digital Equipment Corporation, 8(2):96–116, October 1996. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJM08/DTJM08AH.HTM;   http:
//www.digital.com:80/info/
DTJM08/DTJM08HM.HTM;   http:
//www.digital.com:80/info/
DTJM08/DTJM08P8.PS;   http:

//www.digital.com:80/info/
DTJM08/DTJM08PF.PDF;   http:
//www.digital.com:80/info/
DTJM08/DTJM08SC.TXT.

Lichtenberg:1992:DTA

[LC92] Mitchell P. Lichtenberg and Jeffrey R. Curless. DECnet transport architecture. Digital Technical Journal of Digital Equipment Corporation, 4(1):40–47 (or 40– 46??), Winter 1992. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v4n1/DECnet_Transport_Architecture_
01oct1992DTJ505P8.ps;   http:
//www.digital.com:80/info/
DTJ505/DTJ505SC.TXT.

Leahy:1991:NAF

[Lea91] Lee Leahy. New availability features of local area VAXcluster systems. Digital Technical Journal of Digital Equipment Corporation, 3(3):27– 35, Summer 1991. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v3n3/New_Availability_Features_
of_L_01apr1992DTJ303P8.ps;
http://www.digital.com:
80/info/DTJ303/DTJ303SC.
TXT.

Levine:1994:VRK

[Lev94] Ronald D. Levine. Volume rendering with the Kubota 3D imaging and graphics accelerator. Digital Technical Journal of Digital Equipment Corporation, 6(2):34– 48, Spring 1994. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v6n2/Volume_Rendering_with_
the_Kubo_01oct1994DTJE03P8.
ps;   http://www.digital.
com:80/info/DTJE03/DTJE03SC.
TXT.

Lipchak:1998:PHG

[LFB^{+}98] Benjamin N. Lipchak, Thomas Frisinger, Karen L. Bircsak, Keith L. Comeford, and Michael I. Rosenblum. PowerStorm 4DT: A highperformance graphics software architecture. Digital Technical Journal of Digital Equipment Corporation, 9(4):49–60, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJS04/DTJS04AH.HTM;   http:
//www.digital.com:80/info/
DTJS04/DTJS04HM.HTM;   http:
//www.digital.com:80/info/
DTJS04/DTJS04P8.PS;   http:
//www.digital.com:80/info/
DTJS04/DTJS04PF.PDF;   http:
//www.digital.com:80/info/
DTJS04/DTJS04SC.TXT.

Landau:1991:DDC

[LG91] Richard Landau and Alan Guenther. Design of the DECprint Common Printer Supervisor for VMS system. Digital Technical Journal of Digital Equipment Corporation, 3(4):43–54, Fall 1991. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/

info/DTJ/v3n4/Design_of_
the_DECprint_Common_01jul1992DTJ404P8.
ps;   http://www.digital.
com:80/info/DTJ404/DTJ404SC.
TXT.

Lidington:1988:OVP

[Lid88] Gary P. Lidington. Overview of the MicroVAX 3500/3600 processor module. Digital Technical Journal of Digital Equipment Corporation, 1(7): 79–86, August 1988. CODEN DTJOEL. ISSN 0898-901X.

Laing:1991:TMS

[LJL91] William A. Laing, James E.

Johnson, and Robert V. Landau. Transaction management support in the VMS operating system kernal. Digital Technical Journal of Digital Equipment Corporation, 3(1): 33–44, Winter 1991. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v3n1/Transaction_Management_
Suppor_01oct1991DTJ103P8.
ps;   http://www.digital.
com:80/info/DTJ103/DTJ103SC.
TXT.

LaFrance-Linden:1998:CDH

[LL98] David C. P. LaFrance-Linden.

Challenges in designing an HPF debugger. Digital Technical Journal of Digital Equipment Corporation, 9(3):50–??, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJR04/DTJR04AH.HTM;   http:
//www.digital.com:80/info/

                                        DTJR04/DTJR04HM.HTM;   http:
                                        //www.digital.com:80/info/
                                        DTJR04/DTJR04P8.PS;   http:
                                        //www.digital.com:80/info/
                                        DTJR04/DTJR04PF.PDF;   http:
                                        //www.digital.com:80/info/
                                        DTJR04/DTJR04SC.TXT.

Lauck:1986:DNA

[LOP86] Anthony G. Lauck, David R. Oran, and Radia J. Perlman. Digital Network Architecture overview. Digital Technical Journal of Digital Equipment Corporation, 1(3): 10–24, September 1986. CO-DEN DTJOEL. ISSN 0898-901X.

Larson:1995:SER

[LPWH95] Ray R. Larson, Christian Plaunt, Allison G. Woodruff, and Marti Hearst. The Sequoia 2000 electronic repository. Digital Technical Journal of Digital Equipment Corporation, 7 (3):50–65, Fall 1995. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v7n3/The_Sequoia_2000_Electronic_
                                        Re_02jan1996DTJJ04P8.ps;
                                         http://www.digital.com:
                                        80/info/DTJJ04/DTJJ04AH.
                                        HTM;   http://www.digital.
                                        com:80/info/DTJJ04/DTJJ04P8.
                                        PS;   http://www.digital.
                                        com:80/info/DTJJ04/DTJJ04PF.
                                        PDF;   http://www.digital.
                                        com:80/info/DTJJ04/DTJJ04SC.
                                        TXT.

Laurune:1990:DDI

[LT90] William R. Laurune and Robert L. Travis, Jr. The Digital document interchange format. Digital Technical Journal of Digital Equipment Corporation, 2(1):16–27, Winter 1990. CODEN DTJOEL. ISSN 0898-901X.

Lupton:1988:LE

[Lup88] G. Lupton. Language-sensitive editor. Digital Technical Journal of Digital Equipment Corporation, 1(6):28–39, February 1988. CODEN DTJOEL. ISSN 0898-901X.

Marcello:1996:F

[Mar96] Rich Marcello. Foreword. Digital Technical Journal of Digital Equipment Corporation, 8 (2):??, October 1996. CODEN DTJOEL. ISSN 0898-901X.

URL http://www.digital.
com:80/DTJM00/v8n2-fore.
htm.

Maskas:1986:DMC

[Mas86] Barry A. Maskas. Developing the MicroVAX II CPU board. Digital Technical Journal of Digital Equipment Corporation, 1(2):37–47, March 1986. CODEN DTJOEL. ISSN 0898-901X.

Maskas:1988:DCQ

[Mas88] Barry A. Maskas. Development of the CVAX Q22-bus interface chip. Digital Technical Journal of Digital Equipment Corporation, 1(7):129–138, August 1988. CODEN DTJOEL. ISSN 0898-901X.

Mirchandani:1990:EPR

[MB90] D. Mirchandani and P. Biswas. Ethernet performance of remote DECwindows applications. Digital Technical Journal of Digital Equipment Corporation, 2(3):84–94, Summer 1990. CODEN DTJOEL. ISSN 0898-901X.

McKinney:1994:DDF

[MBC^{+}94] Dina L. McKinney, Masooma Bhaiwala, Kwong-Tak A. Chui, Christopher L. Houghton, James R. Mullens, Daniel L. Leibholz, Sanjay J. Patel, Delvan A. Ramey, and Mark B. Rosenbluth. Digital’s DECchip 21066: The first costfocused Alpha AXP chip. Digital Technical Journal of Digital Equipment Corporation, 6(1): 66–77, Winter 1994. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v6n1/Digitals_DECchip_21066_
The_F_01jul1994DTJD05P8.
ps;   http://www.digital.
com:80/info/DTJD05/DTJD05SC.
TXT.

McElroy:1985:PVP

[McE85] James B. McElroy. Packaging the VAX 8600 processor. Digital Technical Journal of Digital Equipment Corporation, 1 (1):54–60, August 1985. CO-DEN DTJOEL. ISSN 0898-901X.

McGregor:1990:ODA

[McG90] S. A. McGregor. An overview of the DECwindows architecture. Digital Technical Journal of Digital Equipment Corporation, 2(3):9–15, Summer 1990. CODEN DTJOEL. ISSN 0898-901X.

McKeeman:1998:DTS

[McK98] William M. McKeeman. Differential testing for software. Digital Technical Journal of Digital Equipment Corporation, 10(1):100–107, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJT08/DTJT08HM.HTM;   http:
//www.digital.com:80/info/
DTJT08/DTJT08P8.PS;   http:
//www.digital.com:80/info/
DTJT08/DTJT08PF.PDF;   http:
//www.digital.com:80/info/
DTJT08/DTJT08SC.TXT.

Moses:1988:PEV

[MD88] Bhagyam Moses and Karen T.

DeGregory. Performance evaluation of the VAX 6200 systems. Digital Technical Journal of Digital Equipment Corporation, 1(7):64–78, August 1988. CODEN DTJOEL. ISSN 0898-901X.

Methot:1992:CMP

[Met92] Christopher E. Methot. Capacity modeling of PATHWORKS client-server workloads. Digital Technical Journal of Digital Equipment Corporation, 4(1): 68–76, Winter 1992. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/

                                        v4n1/Capacity_Modeling_of_
                                        PATHWORK_01oct1992DTJ508P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJ508/DTJ508SC.
                                        TXT.

Murray:1990:VII

[MHS90] J. E. Murray, R. C. Hetherington, and R. M. Salett. VAX instructions that illustrate the architectural features of the VAX 9000 CPU. Digital Technical Journal of Digital Equipment Corporation, 2(4):25–42, Fall 1990. CODEN DTJOEL. ISSN 0898-901X.

Mierswa:1989:EM

[Mie89] Peter O. Mierswa. The evolution of the MAILbus. Digital Technical Journal of Digital Equipment Corporation, 1 (9):37–43, June 1989. CODEN DTJOEL. ISSN 0898-901X.

Mishra:1987:VM

[Mis87] S. Mishra. The VAX 8800 microarchitecture. Digital Technical Journal of Digital Equipment Corporation, 1(4):20– 33, February 1987. CODEN DTJOEL. ISSN 0898-901X.

Morse:1986:EIE

[MK86] Kathleen D. Morse and Lawrence J. Kenah. The evolution of instruction emulation for the MicroVAX systems. Digital Technical Journal of Digital Equipment Corporation, 1(2): 76–85, March 1986. CODEN DTJOEL. ISSN 0898-901X.

Martin:1989:DVD

[MMO89] Sally J. Martin, Janet M. Mc-Cann, and David R. Oran. Development of the VAX distributed name service. Digital Technical Journal of Digital Equipment Corporation, 1 (9):9–15, June 1989. CODEN DTJOEL. ISSN 0898-901X.

Mierswa:1986:DS

[MMS86] Peter O. Mierswa, David J.

Mitton, and Martha L. Spence. The DECnet-DOS system. Digital Technical Journal of Digital Equipment Corporation, 1(3):108–116, September 1986. CODEN DTJOEL. ISSN 0898-901X.

Morgan:1988:CMCTL

[Mor88] David K. Morgan. The CVAX CMCTL – A CMOS memory controller chip. Digital Technical Journal of Digital Equipment Corporation, 1(7):139– 143, August 1988. CODEN DTJOEL. ISSN 0898-901X.

Morse:1993:F

[Mor93] John A. Morse. Foreword. Digital Technical Journal of Digital Equipment Corporation, 5 (2):??, Spring 1993. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
mm-02-foreword.txt.

Morgan:1998:I

[Mor98] C. Robert Morgan. Introduction. Digital Technical Journal of Digital Equipment Corporation, 10(1):??, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJT00/v10n1-intro.htm.

Morency:1989:MAD

[MPJK89] John P. Morency, Richard P. Pitkin, Ramasamy Jesuraj, and Ambrose C. Kwong. Modeling and analysis of the DECnet/SNA gateway. Digital Technical Journal of Digital Equipment Corporation, 1(9): 87–99, June 1989. CODEN DTJOEL. ISSN 0898-901X.

Morency:1986:DSG

[MPPO86] John P. Morency, David Porter, Richard M. Pitkin, and David R. Oran. The DECnet/SNA gateway product — A case study in cross vendor networking. Digital Technical Journal of Digital Equipment Corporation, 1(3):35–53, September 1986. CODEN DTJOEL. ISSN 0898-901X.

Mann:1986:TSE

[MSK86] Bruce E. Mann, Colin Strutt, and Mark F. Kempf. Terminal servers on Ethernet local area networks. Digital Technical Journal of Digital Equipment Corporation, 1(3): 73–87, September 1986. CO-DEN DTJOEL. ISSN 0898-901X.

Maskas:1992:DPD

[MSW92] Barry A. Maskas, Stephen F. Shirron, and Nicholas A. Warchol. Design and performance of the DEC 4000 AXP departmental server computing systems. Digital Technical Journal of Digital Equipment Corporation, 4(4):82–99, Fall 1992. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v4n4/Design_and_
Performance_of_the_01apr1993DTJ805P8.
ps;   http://www.digital.
com:80/info/DTJ805/DTJ805SC.
TXT.

Montanaro:1997:CRM

[MWA^{+}97] James Montanaro, Richard T.

Witek, Krishna Anne, Andrew J. Black, Elizabeth M. Cooper, Daniel W. Dobberpuhl, Paul M. Donahue, Jim Eno, Gregory W. Hoeppner, David Kruckemyer, Thomas H. Lee, Peter C. M. Lin, Liam Madden, Daniel Murray, Mark H. Pearce, Sribalan Santhanam, Kathryn J. Snyder, Ray Stephany, and Stephen C. Thierauf. A 160-MHz, 32b, 0.5-W CMOS RISC microprocessor. Digital Technical Journal of Digital Equipment Corporation, 9(1):49–??, ???? 1997. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJP05/DTJP05AH.HTM;   http:
//www.digital.com:80/info/
DTJP05/DTJP05HM.HTM;   http:
//www.digital.com:80/info/
DTJP05/DTJP05P8.PS;   http:
//www.digital.com:80/info/
DTJP05/DTJP05PF.PDF;   http:
//www.digital.com:80/info/
DTJP05/DTJP05SC.TXT.

McLellan:1988:DCF

[MWY88] Edward J. McLellan, Gilbert M. Wolrich, and Robert AJ Yodlowski. Development of the CVAX floating point chip. Digital Technical Journal of Digital Equipment Corporation, 1 (7):109–120, August 1988. CO-DEN DTJOEL. ISSN 0898-901X.

Nadkarni:1994:DDP

[NAC^{+}94] Samyojita A. Nadkarni, Walker Anderson, Lauren M. Carlson, David Kravitz, Mitchell O. Norcross, and Thomas M. Wenners. Development of Digital’s PCI chip sets and evaluation kit for the DECchip 21064 microprocessor. Digital Technical Journal of Digital Equipment Corporation, 6(2): 49–61, Spring 1994. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v6n2/Development_of_Digitals_
                                        PCI_C_01oct1994DTJE04P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJE04/DTJE04SC.
                                        TXT.

Nichols:1989:DIV

[NE89] William G. Nichols and Joel S. Emer. Design and implementation of the VAX distributed file service. Digital Technical Journal of Digital Equipment Corporation, 1(9):16–28, June 1989. CODEN DTJOEL. ISSN 0898-901X.

Nemeth:1996:F

[Nem96] Alan G. Nemeth. Foreword.

Digital Technical Journal of Digital Equipment Corporation, 8(3):??, ???? 1996. CO-DEN DTJOEL. ISSN 0898-

901X.  URL http://www.
digital.com/info/DTJN00/
v8n3-fore.htm.

Newcomer:1995:MIA

[New95] Eric A. Newcomer. Multivendor integration architecture: Standards, compliance testing, and applications. Digital Technical Journal of Digital Equipment Corporation, 7(2): 34–46, Spring 1995. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v7n2/Multivendor_Integration_
Archit_02oct1995DTJI03P8.
ps;   http://www.digital.
com:80/info/DTJI03/DTJI03AH.
HTM;   http://www.digital.
com:80/info/DTJI03/DTJI03P8.
PS;   http://www.digital.
com:80/info/DTJI03/DTJI03PF.
PDF;   http://www.digital.
com:80/info/DTJI03/DTJI03SC.
TXT.

Nasr:1992:CTF

[NGBJ92] Andre I. Nasr, Gregory J.

Grula, Antonio C. Berti, and R. D. Jones. CMOS-4 technology for fast logic and dense on-chip memory. Digital Technical Journal of Digital Equipment Corporation, 4(2):39– 50, Spring 1992. CODEN DTJOEL. ISSN 0898-901X.

Notel:1998:OAV

[NK98] Karen L. Notel and Nitin Y. Karkhanis. OpenVMS Alpha 64-bit very large memory design. Digital Technical Journal of Digital Equipment Corporation, 9(4):33–48, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

                                        /www.digital.com:80/info/
                                        DTJS03/DTJS03AH.HTM;   http:
                                        //www.digital.com:80/info/
                                        DTJS03/DTJS03HM.HTM;   http:
                                        //www.digital.com:80/info/
                                        DTJS03/DTJS03P8.PS;   http:
                                        //www.digital.com:80/info/
                                        DTJS03/DTJS03PF.PDF;   http:
                                        //www.digital.com:80/info/
                                        DTJS03/DTJS03SC.TXT.

Neidecker-Lutz:1993:SMP

[NLU93] Burkhard K. Neidecker-Lutz and Robert Ulichney. Software motion pictures. Digital Technical Journal of Digital Equipment Corporation, 5(2): 19–27, Spring 1993. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v5n2/Software_Motion_Pictures_
                                        01oct1993DTJA02P8.ps;   http:
                                        //www.digital.com:80/info/
                                        DTJA02/DTJA02SC.TXT.

Nasr:1992:IDR

[NM92] Mary Beth Nasr and Ellen J. Mager. Implementation of defect reduction strategies into VLSI manufacturing. Digital Technical Journal of Digital Equipment Corporation, 4(2):

73–82, Spring 1992. CODEN DTJOEL. ISSN 0898-901X.

Nourse:1992:MWN

[Nou92]  Andrew W. Nourse.  Mi-

crosoft Windows network virtual device drivers in PATH-WORKS for DOS. Digital Technical Journal of Digital Equipment Corporation, 4(1): 47–55 (or 47–56??), Winter 1992. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v4n1/Microsoft_
Windows_Network_Virt_01oct1992DTJ506P8.
ps;   http://www.digital.
com:80/info/DTJ506/DTJ506SC.
TXT.

Olson:1996:IMD

[OHS96] Margaret Olson, Laura E.

Holly, and Colin Strutt. Integrating multiple directory services. Digital Technical Journal of Digital Equipment Corporation, 8(1):46– 58, May 1996. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v8n1/Integrating_Multiple_
Director_04jun1996DTJL04P8.
ps;   http://www.digital.
com:80/info/DTJL04/DTJL04.
HTM;   http://www.digital.
com:80/info/DTJL04/DTJL04AH.
HTM;   http://www.digital.
com:80/info/DTJL04/DTJL04P8.
PS;   http://www.digital.
com:80/info/DTJL04/DTJL04PF.
PDF;   http://www.digital.
com:80/info/DTJL04/DTJL04SC.
TXT.

Pasquale:1995:HPN

[PAFK95] Joseph Pasquale, Eric W. Anderson, Kevin Fall, and Jonathan S. Kay. Highperformance I/O and networking software in Sequoia 2000. Digital Technical Journal of Digital Equipment Corporation, 7(3):84– 96, Fall 1995. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v7n3/Highperformance_IO_
                                        and_Netwo_02jan1996DTJJ06P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJJ06/DTJJ06AH.
                                        HTM;   http://www.digital.
                                        com:80/info/DTJJ06/DTJJ06P8.
                                        PS;   http://www.digital.
                                        com:80/info/DTJJ06/DTJJ06PF.
                                        PDF;   http://www.digital.
                                        com:80/info/DTJJ06/DTJJ06SC.
                                        TXT.

Pan:1993:DAC

[Pan93] Davis Yen Pan. Digital audio compression. Digital Technical Journal of Digital Equipment Corporation, 5(2):28–33 (or 28– 40??), Spring 1993. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v5n2/Digital_Audio_Compression_
                                        01oct1993DTJA03P8.ps;   http:
                                        //www.digital.com:80/info/
                                        DTJA03/DTJA03SC.TXT.

Patrick:1993:CIU

[Pat93] Paul B. Patrick, Sr. CASE integration using ACA services. Digital Technical Journal of Digital Equipment Corporation, 5(2):84– 99, Spring 1993. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v5n2/CASE_Integration_Using_
ACA_Ser_01oct1993DTJA08P8.
ps;   http://www.digital.
com:80/info/DTJA08/DTJA08SC.
TXT.

Patel:1995:F

[Pat95] Mahendra R. Patel. Foreword. Digital Technical Journal of Digital Equipment Corporation, 7(2):??, Spring 1995. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v7n2_01_02_intro_
fore.txt.

Parodi:1997:IOD

[PB97] J. H. Parodi and F. W.

Burgher. Integrating ObjectBroker and DCE security. Digital Technical Journal of Digital Equipment Corporation, 9(1):42–??, ???? 1997. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJP04/DTJP04AH.HTM;   http:
//www.digital.com:80/info/
DTJP04/DTJP04HM.HTM;   http:
//www.digital.com:80/info/
DTJP04/DTJP04P8.PS;   http:
//www.digital.com:80/info/
DTJP04/DTJP04PF.PDF;   http:
//www.digital.com:80/info/
DTJP04/DTJP04SC.TXT.

Payson:1991:HAB

[PCCW91] Christopher J. Payson, Christopher J. Cianciolo, Robert N. Crouse, and Catherine F. Winsor. Hardware accelerators for bitonal image processing. Digital Technical Journal of Digital Equipment Corporation, 3(4):9–25 (or 9– 26??), Fall 1991. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v3n4/Hardware_Accelerators_
                                        for_Bito_01jul1992DTJ401P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJ401/DTJ401SC.
                                        TXT.

Perlman:1993:RA

[PCS93] Radia J. Perlman, Ross W. Callon, and I. Michael C. Shand. Routing architecture. Digital Technical Journal of Digital Equipment Corporation, 5(1):62– 69, Winter 1993. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v5n1/Routing_Architecture_
                                        01jul1993DTJ905P8.ps;   http:
                                        //www.digital.com:80/info/
                                        DTJ905/DTJ905SC.TXT.

Pesquet:1993:F

[Pes93] Claude Henri Pesquet. Foreword. Digital Technical Journal of Digital Equipment Corporation, 5(3):??, Summer 1993. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.

com/pub/Digital/info/DTJ/
I18N-foreword.txt.

Park:1987:SLP

[PEWQ87] Daeil Park, Rekha D. Von Ehren, Tzyh-Long Wang, and Nii N. Quaynor. System level performance of VAX 8974 and 8978 systems. Digital Technical Journal of Digital Equipment Corporation, 1(5):80–92, September 1987. CODEN DTJOEL. ISSN 0898-901X.

Pledereder:1995:DIO

[PKGV95] Richard Pledereder, Vishu Krishnamurthy, Mayank Gagnon, and M. Vadodaria. DB Integrator: Open middleware for data access. Digital Technical Journal of Digital Equipment Corporation, 7(1): 7–22, Winter 1995. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v7n1/DB_Integrator_Open_
Middlewar_01jul1995DTJH01P8.
ps;   http://www.digital.
com:80/info/DTJH01/DTJH01SC.
TXT.

Porter:1992:VEH

[Por92] Brian Porter. VAX 6000 error handling: a pragmatic approach. Digital Technical Journal of Digital Equipment Corporation, 4(3):92– 104, Summer 1992. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v4n3/VAX_6000_Error_Handling_
A_01jan1993DTJ708P8.ps;

                                        http://www.digital.com:80/
                                        info/DTJ708/DTJ708SC.TXT.

Palmer:1993:DND

[PP93] Lawrence G. Palmer and Ricky S. Palmer. DECspin: A networked desktop videoconferencing application. Digital Technical Journal of Digital Equipment Corporation, 5(2): 65–76, Spring 1993. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v5n2/DECspin_A_Networked_
                                        Desktop_V_01oct1993DTJA06P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJA06/DTJA06SC.
                                        TXT.

Palmer:1998:SDC

[PP98] Lawrence G. Palmer and Ricky S. Palmer. Shared desktop: A collaborative tool for sharing 3-D applications among different window systems. Digital Technical Journal of Digital Equipment Corporation, 9(3):42–??, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

                                        /www.digital.com:80/info/
                                        DTJR03/DTJR03AH.HTM;   http:
                                        //www.digital.com:80/info/
                                        DTJR03/DTJR03HM.HTM;   http:
                                        //www.digital.com:80/info/
                                        DTJR03/DTJR03P8.PS;   http:
                                        //www.digital.com:80/info/
                                        DTJR03/DTJR03PF.PDF;   http:
                                        //www.digital.com:80/info/
                                        DTJR03/DTJR03SC.TXT.

Palmer:1996:PDA

[PR96] Michael Palmer and Jeffrey M.

Russo. Parasight: Debugging and analyzing realtime applications under Digital UNIX. Digital Technical Journal of Digital Equipment Corporation, 7(4):101– 108, March 1996. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v7n4/Parasight_Debugging_
and_Analy_20apr1996DTJK07P8.
ps;   http://www.digital.
com:80/info/DTJK07;   http:
//www.digital.com:80/info/
DTJK07/DTJK07AH.HTM;   http:
//www.digital.com:80/info/
DTJK07/DTJK07P8.PS;   http:
//www.digital.com:80/info/
DTJK07/DTJK07PF.PDF;   http:
//www.digital.com:80/info/
DTJK07/DTJK07SC.TXT.

Pelle:1986:ENM

[PSS86] Nancy R. La Pelle, Mark J.

Seger, and Mark W. Sylor. The evolution of network management products. Digital Technical Journal of Digital Equipment Corporation, 1(3):117– 128, September 1986. CODEN DTJOEL. ISSN 0898-901X.

Piantedosi:1996:PMT

[PSS96] Judith A. Piantedosi, Archana S.

Sathaye, and D. John Shakshober. Performance measurement of TruCluster systems under the TPC-C benchmark. Digital Technical Journal of Digital Equipment Corporation, 8(3):46– 57, December 1996. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v8n3/Performance_Measurement_
                                        of_Tru_07jan1997DTJN04P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJN04/DTJN04AH.
                                        HTM;   http://www.digital.
                                        com:80/info/DTJN04/DTJN04HM.
                                        HTM;   http://www.digital.
                                        com:80/info/DTJN04/DTJN04P8.
                                        PS;   http://www.digital.
                                        com:80/info/DTJN04/DTJN04PF.
                                        PDF;   http://www.digital.
                                        com:80/info/DTJN04/DTJN04SC.
                                        TXT.

Rannenberg:1993:XOI

[RB93] Wendy Rannenberg and Jörgen Bettels. The X/Open internationalization model. Digital Technical Journal of Digital Equipment Corporation, 5 (3):32–42, Summer 1993. CO-DEN DTJOEL. ISSN 0898-

                                        901X.  URL ftp://ftp.
                                        digital.com/pub/Digital/
                                        info/DTJ/v5n3/THE_XOPEN_
                                        INTERNATIONALIZATIO_01jan1994DTJB03P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJB03/DTJB03SC.
                                        TXT.

Rizzolo:1992:DPU

[RBC92] Anthony J. Rizzolo, Elizabeth A. Brewer, and Martha A. Chandler. Design of the PATHWORKS for ULTRIX file server. Digital Technical Journal of Digital Equipment Corporation, 4(1):31–39 (or 31–40??), Winter 1992. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

                                        digital.com/pub/Digital/

info/DTJ/v4n1/Design_of_
the_PATHWORKS_for_01oct1992DTJ504P8.
ps;   http://www.digital.
com:80/info/DTJ504/DTJ504SC.
TXT.

Rengarajan:1996:VCS

[RBG^{+}96] T. K. Rengarajan, Maxwell Berenson, Ganesan Gopal, Bruce McCready, Sapan Panigrahi, Srikant Subramaniam, and Marc B. Sugiyama. VLM capabilities of the Sybase System 11 SQL server. Digital Technical Journal of Digital Equipment Corporation, 8(4):83–??, ???? 1996. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJO07/DTJO07AH.HTM;   http:
//www.digital.com:80/info/
DTJO07/DTJO07HM.HTM;   http:
//www.digital.com:80/info/
DTJO07/DTJO07P8.PS;   http:
//www.digital.com:80/info/
DTJO07/DTJO07PF.PDF;   http:
//www.digital.com:80/info/
DTJO07/DTJO07SC.TXT.

Cao:1987:CBA

[rCQO87] Xi ren Cao, Nii N. Quaynor, and Fernando C. Colon Osorio. CI bus arbitration performance in a VAXcluster system. Digital Technical Journal of Digital Equipment Corporation, 1(5):93–103, September 1987. CODEN DTJOEL. ISSN 0898-901X.

Rege:1991:AIH

[Reg91]  Satish L. Rege.  The ar-

chitecture and implementation of a high-performance FDDI adapter. Digital Technical Journal of Digital Equipment Corporation, 3(3):48– 63, Summer 1991. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v3n3/The_Architecture_and_
Implement_01apr1992DTJ305P8.
ps;   http://www.digital.
com:80/info/DTJ305/DTJ305SC.
TXT.

Reinig:1998:AAD

[Rei98] August G. Reinig. Alias analysis in the DEC C and DIGITAL C++ compilers. Digital Technical Journal of Digital Equipment Corporation, 10(1):48–57, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJT04/DTJT04HM.HTM;   http:
//www.digital.com:80/info/
DTJT04/DTJT04P8.PS;   http:
//www.digital.com:80/info/
DTJT04/DTJT04PF.PDF;   http:
//www.digital.com:80/info/
DTJT04/DTJT04SC.TXT.

Riley:1993:DMO

[RFJR93] Mark F. Riley, James J. Feenan, Jr., John L. Janosik, Jr., and T. K. Rengarajan. The design of multimedia object support in DEC Rdb. Digital Technical Journal of Digital Equipment Corporation, 5 (2):50–64 (or 50–65??), Spring 1993. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v5n2/The_Design_

of_Multimedia_Objec_01oct1993DTJA05P8.
ps;   http://www.digital.
com:80/info/DTJA05/DTJA05SC.
TXT.

Rost:1990:PNT

[RFN90] R. J. Rost, J. D. Friedberg, and P. L. Nishimoto. PEX: a network-transparent three-dimensional graphics system. Digital Technical Journal of Digital Equipment Corporation, 2(3):52–63, Summer 1990. CODEN DTJOEL. ISSN 0898-901X.

Robinson:1993:IOU

[RFW93] David C. Robinson, Lawrence N.

Friedman, and Scott A. Wattum. An implementation of the OSI upper layers and applications. Digital Technical Journal of Digital Equipment Corporation, 5(1):107– 116, Winter 1993. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v5n1/An_Implementation_of_
the_OSI_01jul1993DTJ909P8.
ps;   http://www.digital.
com:80/info/DTJ909/DTJ909SC.
TXT.

Rosenbaum:1996:DCD

[RG96] Richard L. Rosenbaum and Stanley I. Goldfarb. Design of the common directory interface for DECnet/OSI. Digital Technical Journal of Digital Equipment Corporation, 8 (1):59–72 (or 59–67??), May 1996. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v8n1/Design_of_
the_Common_Director_04jun1996DTJL05P8.
ps;   http://www.digital.
com:80/info/DTJL05/DTJL05.
HTM;   http://www.digital.
com:80/info/DTJL05/DTJL05AH.
HTM;   http://www.digital.
com:80/info/DTJL05/DTJL05P8.
PS;   http://www.digital.
com:80/info/DTJL05/DTJL05PF.
PDF;   http://www.digital.
com:80/info/DTJL05/DTJL05SC.
TXT.

Rotithor:1998:MAC

[RHD98] Hemant G. Rotithor, Kevin W. Harris, and Mark W. Davis. Measurement and analysis of C and C++ performance. Digital Technical Journal of Digital Equipment Corporation, 10(1):32–47, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJT03/DTJT03HM.HTM;   http:
//www.digital.com:80/info/
DTJT03/DTJT03P8.PS;   http:
//www.digital.com:80/info/
DTJT03/DTJT03PF.PDF;   http:
//www.digital.com:80/info/
DTJT03/DTJT03SC.TXT.

Rozmovits:1996:DUI

[Roz96] Bernard A. Rozmovits. The design of user interfaces for digital speech recognition software. Digital Technical Journal of Digital Equipment Corporation, 8(2):117–126, October 1996. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/

DTJM09/DTJM09AH.HTM;   http:
//www.digital.com:80/info/
DTJM09/DTJM09HM.HTM;   http:
//www.digital.com:80/info/
DTJM09/DTJM09P8.PS;   http:
//www.digital.com:80/info/
DTJM09/DTJM09PF.PDF;   http:
//www.digital.com:80/info/
DTJM09/DTJM09SC.TXT.

Raghavan:1991:DAT

[RR91] Ananth Raghavan and T. K.

Rengarajan. Database availability for transaction processing. Digital Technical Journal of Digital Equipment Corporation, 3(1):65– 69, Winter 1991. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v3n1/Database_Availability_
for_Tran_01oct1991DTJ106P8.
ps;   http://www.digital.
com:80/info/DTJ106/DTJ106SC.
TXT.

Rengarajan:1989:HAM

[RSW89] T. Rengarajan, P. Spiro, and W. Wright. High availability mechanisms of VAX DBMS software. Digital Technical Journal of Digital Equipment Corporation, 1(8):88– 98, February 1989. CODEN DTJOEL. ISSN 0898-901X.

Roden:1993:FRN

[RT93] Robert J. Roden and Deborah Tayler. Frame relay networks. Digital Technical Journal of Digital Equipment Corporation, 5(1):99– 106, Winter 1993. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v5n1/Frame_Relay_Networks_
                                        01jul1993DTJ908P8.ps;   http:
                                        //www.digital.com:80/info/
                                        DTJ908/DTJ908SC.TXT.

Russo:1994:AS

[Rus94]  Andrew P. Russo.  The

AlphaServer 2100 I/O subsystem. Digital Technical Journal of Digital Equipment Corporation, 6(3):20– 28, Summer 1994. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v6n3/The_AlphaServer_2100_
                                        IO_Subsy_01jan1995DTJF02P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJF02/DTJF02SC.
                                        TXT.

Ryan:1990:DDV

[RV90] M. R. Ryan and J. H. VanGilder. The development of DECwindows VMS mail. Digital Technical Journal of Digital Equipment Corporation, 2(3): 74–83, Summer 1990. CODEN DTJOEL. ISSN 0898-901X.

Schumann:1997:DMC

[Sch97] Reinhard C. Schumann. Design of the 21174 Memory Controller for DIGITAL Personal Workstations. Digital Technical Journal of Digital Equipment Corporation, 9(2):57–??, ???? 1997. CODEN DTJOEL. ISSN 0898-901X. URL http:/

                                        /www.digital.com:80/info/
                                        DTJQ05/DTJQ05AH.HTM;   http:

//www.digital.com:80/info/
DTJQ05/DTJQ05HM.HTM;   http:
//www.digital.com:80/info/
DTJQ05/DTJQ05P8.PS;   http:
//www.digital.com:80/info/
DTJQ05/DTJQ05PF.PDF;   http:
//www.digital.com:80/info/
DTJQ05/DTJQ05SC.TXT.

Sweany:1998:COS

[SCH98] Philip H. Sweany, Steven M.

Carr, and Brett L. Huber. Compiler optimization for superscalar systems: Global instruction scheduling without copies. Digital Technical Journal of Digital Equipment Corporation, 10(1):58–70, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJT05/DTJT05HM.HTM;   http:
//www.digital.com:80/info/
DTJT05/DTJT05P8.PS;   http:
//www.digital.com:80/info/
DTJT05/DTJT05PF.PDF;   http:
//www.digital.com:80/info/
DTJT05/DTJT05SC.TXT.

Sites:1992:BT

[SCK^{+}92] Richard L. Sites, Anton Chernoff, Matthew B. Kirk, Maurice P. Marks, and Scott G. Robinson. Binary translation. Digital Technical Journal of Digital Equipment Corporation, 4(4):137– 152, Fall 1992. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v4n4/Binary_Translation_
01apr1993DTJ809P8.ps;   http:

                                        //www.digital.com:80/info/
                                        DTJ809/DTJ809SC.TXT.

Sullivan:1990:VMS

[SCL^{+}90] Patrick Sullivan, Michael A. Callander, Sr., James R. Lundberg, Rebecca L. Stamm, and William J. Bowhill. The VAX 6000 Model 400 scalar processor module. Digital Technical Journal of Digital Equipment Corporation, 2(2): 27–35, Spring 1990. CODEN DTJOEL. ISSN 0898-901X.

Sinkewicz:1991:UFD

[SCP^{+}91] Ursula Sinkewicz, Chran-Ham Chang, Lawrence G. Palmer, Craig Smelser, and Fred L. Templin. ULTRIX fiber distributed data interface networking subsystem implementation. Digital Technical Journal of Digital Equipment Corporation, 3(2):85–93 (or 85– 94??), Spring 1991. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v3n2/ULTRIX_Fiber_Distributed_
                                        Data_01jan1992DTJ208P8.ps;
                                         http://www.digital.com:
                                        80/info/DTJ208/DTJ208SC.
                                        TXT.

Smith:1997:PPN

[SDLT97] J. Mark Smith, Sean G. Doherty, Oliver J. Leahy, and Dermot M. Tynan. Protecting a private network: The AltaVista firewall. Digital Technical Journal of Digital Equipment Corporation, 9(2):17–??, ???? 1997. CODEN DTJOEL.

ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJQ02/DTJQ02AH.HTM;   http:
//www.digital.com:80/info/
DTJQ02/DTJQ02HM.HTM;   http:
//www.digital.com:80/info/
DTJQ02/DTJQ02P8.PS;   http:
//www.digital.com:80/info/
DTJQ02/DTJQ02PF.PDF;   http:
//www.digital.com:80/info/
DTJQ02/DTJQ02SC.TXT.

Sylor:1993:NM

[SDS93] Mark W. Sylor, Francis Dolan, and David G. Shurtleff. Network management. Digital Technical Journal of Digital Equipment Corporation, 5(1): 117–129, Winter 1993. CO-DEN DTJOEL. ISSN 0898-

901X.  URL ftp://ftp.
digital.com/pub/Digital/
info/DTJ/v5n1/Network_Management_
01jul1993DTJ910P8.ps;   http:
//www.digital.com:80/info/
DTJ910/DTJ910SC.TXT.

Seavey:1992:NDP

[SFK^{+}92] Marden H. Seavey, John V.

Faricelli, Nadim A. Khalil, G. Nanz, L. M. Richardson, C. O. Schiebl, H. R. Soleimani, and M. Thurner. Numerical device and process simulation tools in transistor design. Digital Technical Journal of Digital Equipment Corporation, 4 (2):25–38, Spring 1992. CO-DEN DTJOEL. ISSN 0898-901X.

Slater:1990:VPV

[SFSW90] Debra L. Slater, David M. Fenwick, D. John Shakshober, and Douglas D. Williams. Vector processing on the VAXvector 6000 Model 400. Digital Technical Journal of Digital Equipment Corporation, 2(2): 11–26, Spring 1990. CODEN DTJOEL. ISSN 0898-901X.

Spitz:1986:MMW

[SGZ86] Rick Spitz, Peter George, and Stephen Zalewski. The making of a MicroVAX workstation. Digital Technical Journal of Digital Equipment Corporation, 1(2):56–65, March 1986. CODEN DTJOEL. ISSN 0898-901X.

Shipman:1997:DIS

[Shi97] Nick Shipman. Developing Internet software: AltaVista mail. Digital Technical Journal of Digital Equipment Corporation, 9(2):23–??, ???? 1997. CODEN DTJOEL. ISSN 0898-901X. URL http:/

                                        /www.digital.com:80/info/
                                        DTJQ03/DTJQ03AH.HTM;   http:
                                        //www.digital.com:80/info/
                                        DTJQ03/DTJQ03HM.HTM;   http:
                                        //www.digital.com:80/info/
                                        DTJQ03/DTJQ03P8.PS;   http:
                                        //www.digital.com:80/info/
                                        DTJQ03/DTJQ03PF.PDF;   http:
                                        //www.digital.com:80/info/
                                        DTJQ03/DTJQ03SC.TXT.

Steinman:1996:ACP

[SHK^{+}96] Maurice B. Steinman, George J. Harris, Andrej Kocev, Virginia C. Lamere, and Roger D. Pannell. The AlphaServer 4100 cached processor module architecture and design.

Digital Technical Journal of Digital Equipment Corporation, 8(4):21–??, ???? 1996. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJO02/DTJO02AH.HTM;   http:
//www.digital.com:80/info/
DTJO02/DTJO02HM.HTM;   http:
//www.digital.com:80/info/
DTJO02/DTJO02P8.PS;   http:
//www.digital.com:80/info/
DTJO02/DTJO02PF.PDF;   http:
//www.digital.com:80/info/
DTJO02/DTJO02SC.TXT.

Sichel:1991:ABO

[Sic91] Peter A. Sichel. ACCESS.bus, an open desktop bus. Digital Technical Journal of Digital Equipment Corporation, 3 (4):36–42, Fall 1991. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v3n4/ACCESS.bus_an_Open_
Desktop_Bu_01jul1992DTJ403P8.
ps;   http://www.digital.
com:80/info/DTJ403/DTJ403SC.
TXT.

Sicola:1994:ADH

[Sic94] Stephen J. Sicola. The architecture and design of HSseries StorageWorks Array Controllers. Digital Technical Journal of Digital Equipment Corporation, 6(4):5– 25, Fall 1994. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v6n4/The_Architecture_and_
Design_o_01apr1995DTJG01P8.

                                        ps;   http://www.digital.
                                        com:80/info/DTJG01/DTJG01SC.
                                        TXT.

Sidman:1989:CST

[Sid89] Michael D. Sidman. Control systems technology in digital’s disk drives. Digital Technical Journal of Digital Equipment Corporation, 1(8): 61–73, February 1989. CODEN DTJOEL. ISSN 0898-901X.

Sites:1992:AAA

[Sit92] Richard L. Sites. Alpha AXP architecture. Digital Technical Journal of Digital Equipment Corporation, 4(4):19–34, Fall 1992. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

                                        digital.com/pub/Digital/
                                        info/DTJ/v4n4/Alpha_AXP_
                                        Architecture_01apr1993DTJ801P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJ801/DTJ801SC.
                                        TXT.

Sites:1995:F

[Sit95]  Richard L. Sites.  Fore-

word. Digital Technical Journal of Digital Equipment Corporation, 7(1):5– 6, Winter 1995. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v7n1-00-foreword.txt.html.

Spiro:1991:DOT

[SJR91] Peter M. Spiro, Ashok M. Joshi, and T. K. Rengarajan. Designing an optimized transaction commit protocol. Digital Technical Journal of Digital Equipment Corporation, 3 (1):70–78 (or 70–79??), Winter 1991. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v3n1/Designing_
an_Optimized_Transac_01oct1991DTJ107P8.
ps;   http://www.digital.
com:80/info/DTJ107/DTJ107SC.
TXT.

Sung:1990:DDD

[SJY90] Alan Sung, Neal F. Jacobson, and Carol A. Young. The design and development of the DECdecision product. Digital Technical Journal of Digital Equipment Corporation, 2(1): 60–72, Winter 1990. CODEN DTJOEL. ISSN 0898-901X.

Souza:1994:GSH

[SKÖ^{+}94] Robert J. Souza, P. G. Krishnakumar, Cüneyt M. Özveren, Robert J. Simcoe, Barry A. Spinney, Robert E. Thomas, and Robert J. Walsh. GI-GAswitch system: A highperformance packet-switching platform. Digital Technical Journal of Digital Equipment Corporation, 6(1):9– 22, Winter 1994. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v6n1/GIGAswitch_System_A_
Highperf_01jul1994DTJD01P8.
ps;   http://www.digital.
com:80/info/DTJD01/DTJD01SC.
TXT.

Smith:1989:MDO

[Smi89] Alan B. Smith. Magnetic domain observations in thin-film heads using kerr microscopy. Digital Technical Journal of Digital Equipment Corporation, 1(8):74–80, February 1989. CODEN DTJOEL. ISSN 0898-901X.

Smith:1996:APS

[Smi96] Duane A. Smith. Adding 64bit pointer support to a 32bit run-time library. Digital Technical Journal of Digital Equipment Corporation, 8 (2):83–95 (or 83–93??), October 1996. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJM07/DTJM07AH.HTM;   http:
//www.digital.com:80/info/
DTJM07/DTJM07HM.HTM;   http:
//www.digital.com:80/info/
DTJM07/DTJM07P8.PS;   http:
//www.digital.com:80/info/
DTJM07/DTJM07PF.PDF;   http:
//www.digital.com:80/info/
DTJM07/DTJM07SC.TXT.

Simone:1991:DTP

[SMV91] Guido Simone, Jeffrey A. Metzger, and Gary Vaillette. Design of the Turbo PrintServer 20 controller. Digital Technical Journal of Digital Equipment Corporation, 3(4):61–72 (or 61– 76??), Fall 1991. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v3n4/Design_of_the_Turbo_
PrintServe_01jul1992DTJ406P8.
ps;   http://www.digital.
com:80/info/DTJ406/DTJ406SC.
TXT.

Snaman:1991:ADV

[Sna91] William E. Snaman, Jr. Application design in a VAXcluster system. Digital Technical Journal of Digital Equipment Corporation, 3(3):16– 26, Summer 1991. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v3n3/Application_Design_
in_a_VAXclu_01apr1992DTJ302P8.
ps;   http://www.digital.
com:80/info/DTJ302/DTJ302SC.
TXT.

Speer:1991:DTP

[SS91] Thomas G. Speer and Mark W.

Storm. Digital’s transaction processing monitors. Digital Technical Journal of Digital Equipment Corporation, 3(1): 18–32, Winter 1991. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v3n1/Digitals_Transaction_
Processi_01oct1991DTJ102P8.
ps;   http://www.digital.
com:80/info/DTJ102/DTJ102SC.
TXT.

Strutt:1993:DDM

[SS93] Colin Strutt and James A.

Swist. Design of the DECmcc Management Director. Digital Technical Journal of Digital Equipment Corporation, 5(1):130–142, Winter 1993. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v5n1/Design_of_

                                        the_DECmcc_Managemen_01jul1993DTJ911P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJ911/DTJ911SC.
                                        TXT.

Snaman:1987:VVD

[ST87] William E. Snaman, Jr. and David W. Thiel. The VAX/ VMS distributed lock manager. Digital Technical Journal of Digital Equipment Corporation, 1(5):29–44, September 1987. CODEN DTJOEL. ISSN 0898-901X.

Stonebraker:1995:OSP

[Sto95]  Michael Stonebraker.  An

overview of the Sequoia 2000 project. Digital Technical Journal of Digital Equipment Corporation, 7(3):39– 49, Fall 1995. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v7n3/An_Overview_of_the_
                                        Sequoia_02jan1996DTJJ03P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJJ03/DTJJ03AH.
                                        HTM;   http://www.digital.
                                        com:80/info/DTJJ03/DTJJ03P8.
                                        PS;   http://www.digital.
                                        com:80/info/DTJJ03/DTJJ03PF.
                                        PDF;   http://www.digital.
                                        com:80/info/DTJJ03/DTJJ03SC.
                                        TXT.

Seiler:1996:IVR

[SU96] Larry D. Seiler and Robert A. Ulichney. Integrating video rendering into graphics accelerator chips. Digital Technical Journal of Digital Equipment Corporation, 7(4):76– 88, March 1996. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v7n4/Integrating_Video_Rendering_
in_20apr1996DTJK05P8.ps;
http://www.digital.com:
80/info/DTJK05;   http://
www.digital.com:80/info/
DTJK05/DTJK05AH.HTM;   http:
//www.digital.com:80/info/
DTJK05/DTJK05P8.PS;   http:
//www.digital.com:80/info/
DTJK05/DTJK05PF.PDF;   http:
//www.digital.com:80/info/
DTJK05/DTJK05SC.TXT.

Supnik:1992:Fc

[Sup92a]  Robert M. Supnik.  Fore-

word. Digital Technical Journal of Digital Equipment Corporation, 4(3):9–??, Summer 1992. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/nv-foreword.txt.

Supnik:1992:Fd

[Sup92b]  Robert M. Supnik.  Fore-

word. Digital Technical Journal of Digital Equipment Corporation, 4(4):??, Fall 1992. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/axp-foreword.txt.

Spine:1990:EXU

[SV90] T. M. Spine and J. L. VanNoy.

The evolution of the X user interface style. Digital Technical Journal of Digital Equipment Corporation, 2(3):44–51, Summer 1990. CODEN DTJOEL. ISSN 0898-901X.

Stockdale:1991:DDL

[SW91] Richard E. Stockdale and Judy B. Weiss. Design of the DEC LANcontroller 400 adapter. Digital Technical Journal of Digital Equipment Corporation, 3(3):36– 47, Summer 1991. CO-DEN DTJOEL. ISSN 0898-

901X.  URL ftp://ftp.
digital.com/pub/Digital/
info/DTJ/v3n3/Design_of_
the_DEC_LANcontrolle_01apr1992DTJ304P8.
ps;   http://www.digital.
com:80/info/DTJ304/DTJ304SC.
TXT.

Sweet:1991:DMD

[Swe91] Bruce E. Sweet. DECelms — managing Digital’s FDDI and Ethernet extended local area networks. Digital Technical Journal of Digital Equipment Corporation, 3(2):76– 84, Spring 1991. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v3n2/DECelms_Managing_Digitals_
F_01jan1992DTJ207P8.ps;
http://www.digital.com:80/
info/DTJ207/DTJ207SC.TXT.

Sylor:1986:NDM

[Syl86] Mark W. Sylor. The NMCC/ DECnet monitor design. Digital Technical Journal of Digital Equipment Corporation, 1 (3):129–141, September 1986. CODEN DTJOEL. ISSN 0898-901X.

Tallman:1995:PGA

[Tal95] Owen H. Tallman. Project Gabriel: Automated software deployment in a large commercial network. Digital Technical Journal of Digital Equipment Corporation, 7(2):56– 70, Spring 1995. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v7n2/Project_Gabriel_Automated_
Sof_02oct1995DTJI05P8.ps;
 http://www.digital.com:
80/info/DTJI05/DTJI05AH.
HTM;   http://www.digital.
com:80/info/DTJI05/DTJI05P8.
PS;   http://www.digital.
com:80/info/DTJI05/DTJI05PF.
PDF;   http://www.digital.
com:80/info/DTJI05/DTJI05SC.
TXT.

Treggiari:1990:DXT

[TC90] L. P. Treggiari and M. D.

Collins. Development of the XUI toolkit. Digital Technical Journal of Digital Equipment Corporation, 2(3):24–33, Summer 1990. CODEN DTJOEL. ISSN 0898-901X.

Troiani:1985:VBP

[TCQ^{+}85] Mario Troiani, S. Stephen Ching, Nii N. Quaynor, John E. Bloem, and Fernando C. Colon Osorio. The VAX 8600 I box, A pipelined implementation of the VAX architecture. Digital Technical Journal of Digital Equipment Corporation, 1(1): 24–42, August 1985. CODEN DTJOEL. ISSN 0898-901X.

Thacker:1992:ADU

[TCS92] Charles P. Thacker, David G. Conroy, and Lawrence C. Stewart. The Alpha demonstration unit: A high-performance multiprocessor for software and chip development. Digital Technical Journal of Digital Equipment Corporation, 4 (4):51–65, Fall 1992. CO-DEN DTJOEL. ISSN 0898-

                                        901X.  URL ftp://ftp.
                                        digital.com/pub/Digital/
                                        info/DTJ/v4n4/The_Alpha_
                                        Demonstration_Unit_01apr1993DTJ803P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJ803/DTJ803SC.
                                        TXT.

teKiefte:1993:MID

[tHMv93] Jan B. te Kiefte, Bob Hasenaar, Joop W. Mevius, and Theo M. van Hunnik. The Megadoc image document management system. Digital Technical Journal of Digital Equipment Corporation, 5(2): 41–49, Spring 1993. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v5n2/The_Megadoc_Image_Document_
                                        Man_01oct1993DTJA04P8.ps;
                                         http://www.digital.com:
                                        80/info/DTJA04/DTJA04SC.
                                        TXT.

Thomson:1993:AQO

[Tho93] Robert G. Thomson. Assessing the quality of Open-VMS AXP: Software measurement using subjective data. Digital Technical Journal of Digital Equipment Corporation, 5(4):69–78, Fall 1993. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v5n4/Assessing_
the_Quality_of_OpenV_01apr1994DTJC06P8.
ps;   http://www.digital.
com:80/info/DTJC06/DTJC06SC.
TXT.

Tiffany:1991:DP

[TKK91] William J. Tiffany, G. Paul Koning, and James E. Kuenzel. The DECconcentrator 500 product. Digital Technical Journal of Digital Equipment Corporation, 3(2):64– 75, Spring 1991. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v3n2/The_DECconcentrator_
500_Produc_01jan1992DTJ206P8.
ps;   http://www.digital.
com:80/info/DTJ206/DTJ206SC.
TXT.

Travis:1990:COC

[Tra90] Robert L. Travis, Jr. CDA overview (compound document architecture). Digital Technical Journal of Digital Equipment Corporation, 2(1):8–15, Winter 1990. CODEN DTJOEL. ISSN 0898-901X.

Treggiari:1997:DFM

[Tre97]  L. P. Treggiari.  Develop-

ment of the Fortran Module Wizard within DIGITAL Visual Fortran. Digital Technical Journal of Digital Equipment Corporation, 9(1):13–??, ???? 1997. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJP02/DTJP02AH.HTM;   http:
//www.digital.com:80/info/
DTJP02/DTJP02HM.HTM;   http:
//www.digital.com:80/info/
DTJP02/DTJP02P8.PS;   http:
//www.digital.com:80/info/
DTJP02/DTJP02PF.PDF;   http:
//www.digital.com:80/info/
DTJP02/DTJP02SC.TXT.

Uhler:1992:NNH

[UBB^{+}92] G. Michael Uhler, Debra Bernstein, Larry L. Biro, John F. Brown III, John H. Edmondson, Jeffrey D. Pickholtz, and Rebecca L. Stamm. The NVAX and NVAX+ high-performance VAX microprocessors. Digital Technical Journal of Digital Equipment Corporation, 4(3): 11–23, Summer 1992. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v4n3/The_NVAX_and_NVAX_Highperfor_
01jan1993DTJ701P8.ps;   http:
//www.digital.com:80/info/
DTJ701/DTJ701SC.TXT.

Ulichney:1993:VR

[Uli93]  Robert Ulichney.  Video

rendering. Digital Technical Journal of Digital Equipment Corporation, 5(2):9– 18, Spring 1993. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v5n2/Video_Rendering_01oct1993DTJA01P8.
ps;   http://www.digital.

com:80/info/DTJA01/DTJA01SC.
TXT.

Ulichney:1996:F

[Uli96] Robert Ulichney. Foreword.

Digital Technical Journal of Digital Equipment Corporation, 7(4):??, March 1996. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJK00/v7n4-fore.htm.

Viscarola:1989:WTB

[VW89] Peter G. Viscarola and Jeffrey E. Watkins. The WAVE tools base for protocol testing. Digital Technical Journal of Digital Equipment Corporation, 1(9):51–60, June 1989. CODEN DTJOEL. ISSN 0898-901X.

Walsh:1998:DFA

[Wal98] Robert J. Walsh. DART:

Fast application-level networking via data-copy avoidance. Digital Technical Journal of Digital Equipment Corporation, 9(4):61–75, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJS05/DTJS05AH.HTM;   http:
//www.digital.com:80/info/
DTJS05/DTJS05HM.HTM;   http:
//www.digital.com:80/info/
DTJS05/DTJS05P8.PS;   http:
//www.digital.com:80/info/
DTJS05/DTJS05PF.PDF;   http:
//www.digital.com:80/info/
DTJS05/DTJS05SC.TXT.

Whitaker:1996:DSS

[WBW96] Christian Whitaker, J. Stuart Bayley, and Rod D. W. Widdowson. Design of the server for the Spiralog File System. Digital Technical Journal of Digital Equipment Corporation, 8(2):15–31, October 1996. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJM02/DTJM02AH.HTM;   http:
//www.digital.com:80/info/
DTJM02/DTJM02HM.HTM;   http:
//www.digital.com:80/info/
DTJM02/DTJM02P8.PS;   http:
//www.digital.com:80/info/
DTJM02/DTJM02PF.PDF;   http:
//www.digital.com:80/info/
DTJM02/DTJM02SC.TXT.

Weber:1990:COR

[Web90] Larry B. Weber. Compiler optimization in RISC systems. Digital Technical Journal of Digital Equipment Corporation, 2 (2):89–95, Spring 1990. CO-DEN DTJOEL. ISSN 0898-901X.

Wells:1992:DOP

[Wel92] Philip J. Wells. The development of an optimized PATHWORKS transport interface. Digital Technical Journal of Digital Equipment Corporation, 4(1):24–30 (or 24– 31??), Winter 1992. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v4n1/The_Development_of_
an_Optimize_01oct1992DTJ503P8.

ps;   http://www.digital.
com:80/info/DTJ503/DTJ503SC.
TXT.

Weiss:1997:DPW

[WH97] Kenneth M. Weiss and Kenneth A. House. DIGITAL Personal Workstations: The design of high-performance, lowcost Alpha systems. Digital Technical Journal of Digital Equipment Corporation, 9(2):45–??, ???? 1997. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJQ04/DTJQ04AH.HTM;   http:
//www.digital.com:80/info/
DTJQ04/DTJQ04HM.HTM;   http:
//www.digital.com:80/info/
DTJQ04/DTJQ04P8.PS;   http:
//www.digital.com:80/info/
DTJQ04/DTJQ04PF.PDF;   http:
//www.digital.com:80/info/
DTJQ04/DTJQ04SC.TXT.

Winston:1988:SSC

[Win88] Jeff Winston. The system support chip, a multifunction chip for CVAX systems. Digital Technical Journal of Digital Equipment Corporation, 1 (7):121–128, August 1988. CO-DEN DTJOEL. ISSN 0898-901X.

Winters:1993:IDS

[Win93]  Gayn B. Winters.  Inter-

national distributed systems — architectural and practical issues. Digital Technical Journal of Digital Equipment Corporation, 5(3):53– 62, Summer 1993. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v5n3/INTERNATIONAL_DISTRIBUTED_
SYST_01jan1994DTJB05P8.ps;
 http://www.digital.com:
80/info/DTJB05/DTJB05SC.
TXT.

Wilson:1996:DBO

[WNR96] Linda S. Wilson, Craig A. Neth, and Michael J. Rickabaugh. Delivering binary object modification tools for program analysis and optimization. Digital Technical Journal of Digital Equipment Corporation, 8(1):19– 31 (or 18–31??), May 1996. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v8n1/Delivering_
Binary_Object_Modif_04jun1996DTJL02P8.
ps;   http://www.digital.
com:80/info/DTJL02/DTJL02.
HTM;   http://www.digital.
com:80/info/DTJL02/DTJL02AH.
HTM;   http://www.digital.
com:80/info/DTJL02/DTJL02P8.
PS;   http://www.digital.
com:80/info/DTJL02/DTJL02PF.
PDF;   http://www.digital.
com:80/info/DTJL02/DTJL02SC.
TXT.

Warchol:1986:RQDX

[WS86] Nicholas A. Warchol and Stephen F. Shirron. The RQDX3 design project. Digital Technical Journal of Digital Equipment Corporation, 1(2): 66–75, March 1986. CODEN DTJOEL. ISSN 0898-901X.

Yau:1993:SCJ

[Yau93] Michael M. T. Yau. Supporting the Chinese, Japanese, and Korean languages in the OpenVMS operating system. Digital Technical Journal of Digital Equipment Corporation, 5(3):63–79, Summer 1993. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v5n3/SUPPORTING_
THE_CHINESE_JAPANE_01jan1994DTJB06P8.
ps;   http://www.digital.
com:80/info/DTJB06/DTJB06SC.
TXT.

Young:1990:DTI

[YJ90] Carol A. Young and Neal F.

Jacobson. The Digital table interchange format. Digital Technical Journal of Digital Equipment Corporation, 2 (1):28–37, Winter 1990. CO-DEN DTJOEL. ISSN 0898-901X.

Yoshioka:1993:CID

[YM93] Hirotaka Yoshioka and Jim Melton. Character internationalization in databases: A case study. Digital Technical Journal of Digital Equipment Corporation, 5(3):80–96, Summer 1993. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v5n3/CHARACTER_
INTERNATIONALIZATIO_01jan1994DTJB07P8.
ps;   http://www.digital.
com:80/info/DTJB07/DTJB07SC.
TXT.

Yetto:1993:DOO

[YMP^{+}93] Lawrence Yetto, Dorothy Noren Millbrandt, Yanick Pouffary, Daniel J. Ryan Jr., and David J. Sullivan. The DECnet/OSI for OpenVMS version 5.5 implementation. Digital Technical Journal of Digital Equipment Corporation, 5(1):21–33 (or 1–13??), Winter 1993. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.
                                        com/pub/Digital/info/DTJ/
                                        v5n1/The_DECnetOSI_for_OpenVMS_
                                        Ver_01jul1993DTJ902P8.ps;
                                         http://www.digital.com:
                                        80/info/DTJ902/DTJ902SC.
                                        TXT.

Yang:1991:FDL

[YST91] Henry S. Yang, Barry A. Spinney, and Stephen Towning. FDDI data link development. Digital Technical Journal of Digital Equipment Corporation, 3(2):31–41, Spring 1991. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

                                        digital.com/pub/Digital/
                                        info/DTJ/v3n2/FDDI_Data_
                                        Link_Developmentysic_01jan1992DTJ203P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJ203/DTJ203SC.
                                        TXT.

Zaremba:1995:ADV

[Zar95] Donald A. Zaremba. Adding a data visualization tool to DEC FUSE. Digital Technical Journal of Digital Equipment Corporation, 7(2):20– 33, Spring 1995. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v7n2/Adding_a_Data_Visualization_
To_02oct1995DTJI02P8.ps;
 http://www.digital.com:
80/info/DTJI02/DTJI02AH.
HTM;   http://www.digital.
com:80/info/DTJI02/DTJI02P8.
PS;   http://www.digital.
com:80/info/DTJI02/DTJI02PF.
PDF;   http://www.digital.
com:80/info/DTJI02/DTJI02SC.
TXT.

Ziman:1988:PMV

[ZD88] L. Ziman and M. Dickau.

Project management of the VAX DEC/Test Manager software version 2.0. Digital Technical Journal of Digital Equipment Corporation, 1(6):110– 116, February 1988. CODEN DTJOEL. ISSN 0898-901X.

Zetterlund:1992:MPP

[ZFF92] B. Zetterlund, J. A. Farrell, and T. F. Fox. Microprocessor performance and process complexity in CMOS technologies. Digital Technical Journal of Digital Equipment Corporation, 4(2):12–24, Spring 1992. CODEN DTJOEL. ISSN 0898-901X.

Zahavi:1991:TTP

[ZHO91] William Z. Zahavi, Frances A.

Habib, and Kenneth J. Omahen. Tools and techniques for preliminary sizing of transaction processing applications. Digital Technical Journal of Digital Equipment Corporation, 3(1):58–64, Winter 1991. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

                                        digital.com/pub/Digital/
                                        info/DTJ/v3n1/Tools_and_
                                        Techniques_for_Preli_01oct1991DTJ105P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJ105/DTJ105SC.
                                        TXT.

Zurawski:1995:DVA

[ZML95] John H. Zurawski, John E. Murray, and Paul J. Lemmon. The design and verification of the AlphaStation 600 5-series workstation. Digital Technical Journal of Digital Equipment Corporation, 7 (1):89–99, Winter 1995. CO-DEN DTJOEL. ISSN 0898-

                                        901X.  URL ftp://ftp.
                                        digital.com/pub/Digital/
                                        info/DTJ/v7n1/The_Design_
                                        and_Verification_o_01jul1995DTJH07P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJH07/DTJH07SC.
                                        TXT.